### v1

# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

Abstract
Large Language Models, though successful in code generation, struggle with code quality analysis because they are limited by static training data and can’t easily adapt to evolving best practices. We introduce MetaLint, a new instruction-following framework that formulates code quality analysis as the task of detecting and fixing problematic semantic code fragments or code idioms based on high-level specifications. Unlike conventional approaches that train models on static, rule-based data, MetaLint employs instruction tuning on synthetic linter-generated data to support easy-to-hard generalization, enabling models to adapt to novel or complex code patterns without retraining. To evaluate this, we construct a benchmark of challenging idioms inspired by real-world coding standards such as Python Enhancement Proposals (PEPs) and assess whether MetaLint-trained models reason adaptively or simply memorize. Our results show that MetaLint improves generalization to unseen PEP idioms, achieving a 70.37% F-score on idiom detection with the highest recall (70.43%) among all evaluated models. It also achieves 26.73% on localization, competitive for its 4B parameter size and comparable to larger state-of-the-art models like o3-mini, highlighting its potential for future-proof code quality analysis.
1 Introduction
With the rise of Large Language Model (LLM) agents, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent. Researchers have been investigating the potential of LLMs to evaluate and improve code quality through benchmarks (Chambon et al., 2025; Zheng et al., 2024; Waghjale et al., 2024), code review agents (Vijayvergiya et al., 2024; Rasheed et al., 2024), and static analysis with LLMs (Fang et al., 2025; Holden & Kahani, 2024; Khare et al., 2023). However, many of these efforts are hindered by reliance on task-specific training data, often based on narrow or outdated coding best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024). As a result, these systems often perform poorly when detecting rare issue types or when applied to code distributions that differ from their training data (Holden & Kahani, 2024). They may also over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). Ideally, we would develop LLM systems capable of identifying code quality issues without explicit supervision for the target idioms, especially for hard or rare patterns, and that can adapt to evolving best practices over time.
One approach to this problem is to train the LLM on a more general task: understanding and detecting semantic blocks of code, also known as code idioms. For example, a commonly used idiom for generating secrets or passwords in Python is to use the random.choice standard library function. However, as noted in PEP 506 (D’Aprano, 2017), it is cryptographically insecure and Python documentation explicitly warns against the usage of this module for security reasons, which is often missed by developers as highlighted by accepted answers on forums like StackOverflow. PEP 506 also introduces a more secure semantic block or idiom in the form of the secrets module and the secrets.choice function, which acts as a safer alternative to the random.choice idiom. As illustrated by this example, detecting and locating idioms associated with bad practices can be leveraged for identifying code quality issues like code smells (Wikipedia contributors, 2024) or Common Weakness Enumerations (CWE) (MITRE Corporation, 2024). Additionally, the issues can be fixed by replacing instances of “bad” idioms with corresponding “good” idioms that align with best practices.
In this work we extended this observation by training the LLM on a higher-level instruction-following task: given a specification of a best-practice code idiom $I$, the model learns to identify and localize code fragments that violate the idiom and could be rewritten to adhere to or incorporate idiom $I$. We refer to this instruction-following task as meta-linting. The core idea is to teach the LLM the underlying structure of the task, namely identifying and locating idioms based on high-level specifications, rather than training it to memorize or learn a static set of best practices. Since the focus is on improving the model’s instruction-following ability rather than tailoring it to a particular data distribution of best practices, this approach can keep up with evolving best practices.
Our approach, MetaLint, is motivated by the observation that instruction tuning enables cross-task generalization and improves performance on unseen tasks (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Since meta linting frames each idiom as a separate task, instruction style training is naturally suited to generalize to novel and unseen idioms. Existing linters for programming languages (e.g., Ruff for Python) implement numerous simple rules, or “easy” idioms, that reflect aspects of code quality such as modern syntax, readability, and security. These linters can be used to generate large-scale synthetic data for supervised instruction fine-tuning and serve as verifiers during preference optimization or reinforcement learning. This setup also supports easy-to-hard generalization: training on automatically generated linter data for easy idioms helps build a strong model that can handle harder, more abstract idioms at test time, which go beyond the reach of static analyzers. Similar strategies have shown success in math reasoning tasks (Sun et al., 2024b). To systematically study this generalization, we construct a challenging benchmark of hard idioms based on popular PEPs that introduce high-level constructs. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them to models trained with MetaLint, focusing on whether LLMs can move beyond memorization, generalize from easy to hard idioms, and adapt to novel patterns across code quality domains.
Our key contributions are as follows:
  1. 1.
    We introduce MetaLint, an instruction-following framework that leverages synthetic data for easy-to-hard generalization and addresses the limitations of static training used in prior LLM-based code quality approaches.
  2. 2.
    We construct a benchmark of challenging, widely relevant idioms inspired by various PEPs, and evaluate state-of-the-art code and reasoning models on it.
  3. 3.
    We analyze whether MetaLint promotes adaptive reasoning over memorization and enables generalization from easy to hard idioms.
2 Related Work
2.1 Code Quality Analysis with Large Language Models
A large body of prior work has explored the use of LLMs for code quality analysis through code review and static analysis. Tools like GPTLint (Travis Fischer, 2024) and lintrule (lin, 2023) treat LLMs as linters guided by rule descriptions, relying on prompting or fine-tuning to improve rule enforcement. LiteLLM (Fang et al., 2025) leverages LLMs to reduce both the cost and false positive rate of linting for Verilog. Khare et al. (2023) show that LLMs can outperform traditional static analysis tools for detecting certain security-related CWEs, especially when step-by-step reasoning is used. Vijayvergiya et al. (2024) train LLMs to detect and localize violations of best practices, while Rasheed et al. Rasheed et al. (2024) design a multi-agent review pipeline targeting issues like maintainability, efficiency, and bugs. Other works (Jiang et al., 2025b; Yao et al., 2025) use comparative prefix-tuning and reinforcement learning with static analysis–based rewards to promote the generation of higher-quality functionally correct code. Naik et al. (2024) and Jaoua et al. (2025) integrate LLMs with linters to produce more informative code reviews. Finally, CoUpJava (Jiang et al., 2025a) presents a benchmark of Java version–related code upgrade histories, conceptually similar to the hard PEP idiom benchmark we developed for Python. While such efforts highlight the potential of LLMs in code quality tasks, they often target specific rule sets or best practices, necessitating retraining as those evolve. In contrast, our instruction-following meta-linting framework trains models to interpret high-level code quality specifications and perform static analysis to enforce them, enabling broader generalization beyond fixed rules.
2.2 Instruction Following for Generalization
Instruction tuning has emerged as a powerful form of meta-learning that enables cross-task generalization by training models to interpret and follow natural language instructions rather than learning a single fixed task. Prior work shows that exposing models to diverse task instructions allows them to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, acting as a substitute for supervision (Puri et al., 2022) and enabling models to generalize even with minimal labeled data or via pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity is a key driver of generalization, with varied instructions improving performance more than repeated exposure to the same task format (Charton et al., 2024). This phenomenon holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) also performs task-specific instruction following using synthetic data, demonstrating its effectiveness; however, it relies entirely on LLM-generated data and does not incorporate verifiers. Collectively, these results suggest that instruction tuning functions as task-level meta-learning, equipping models with the ability to rapidly adapt to new tasks via natural language. Inspired by this work we treat modeling of specific code quality idioms as individual tasks and propose the MetaLint training framework that generates synthetic data at scale for each meta-task to facilitate cross-idiom generalization. This way the trained model can keep up with new idioms and evolving best practices.
2.3 Easy-to-Hard Generalization
Recent research demonstrates that easy-to-hard generalization improves model performance across mathematical, algorithmic, and coding tasks, motivating its application to code quality analysis. In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025). Several works implement curriculum strategies to structure this progression (Parashar et al., 2025; Zhao et al., 2025), while others emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024). These principles extend beyond math: Sun et al. (2024a) shows that reward models trained on simple code and math problems improve performance on complex ones, and Zhao et al. (2025) further explores generalization in code tasks through curriculum fine-tuning. Multi-task and length generalization studies (Hu et al., 2025) and differentiable programming setups (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code. Zhang et al. (2024) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats. While our approach does not explicitly follow curriculum learning, we draw inspiration from this line of work by training MetaLint on large-scale synthetic data that captures easily detectable code idioms, many of which are directly covered by rule-based linters. We hypothesize that these simple patterns act as a stepping stone, enabling generalization to more complex and abstract unseen PEP idioms.
3 Method
Figure 1: MetaLint training framework (1/2): Core components of our framework include using linter tools (like Ruff) for 1) Synthetic Data Generation, 2) Supervised Instruction Fine-Tuning (SFT) with synthetic data and a 3) Verifiable Reward Model (also derived from the Ruff linter).
Figure 2: MetaLint training framework (2/2): We use the linter-based verifiable reward model for the final two preference optimization components: 4) Rejection Sampling Direct Preference Optimization (RS-DPO), and, 5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
We propose the MetaLint framework to support the learning of skills essential for code idiom analysis, such as interpreting idiom descriptions in terms of underlying code constructs and performing line-by-line static analysis to detect idiom violations. We formulate idiom detection as an instruction-following meta-task $M_{I}$ for a given idiom $I$, where the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}={D_{I},E_{I}}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization and encourages adaptive reasoning over the prompt’s specification, since flagging violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach enables the LLM to remain flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
3.1 Synthetic Data Generation
A key advantage of our meta-task formulation is its support for easy-to-hard generalization. We train LLMs on a set of idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$ to improve their ability to flag idiom violations, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by $\mathcal{L}$). Our hypothesis is that training on $I_{\mathcal{L}}$ helps the LLM acquire the ability to understand, and detect code idioms from high level descriptions, enabling it to generalize more effectively to the harder idioms in $I_{\mathcal{L^{\prime}}}$ compared to an untrained checkpoint.
Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For this purpose, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, semantics, and more. We run Ruff on Python source code files $f\in\mathcal{F}$ from the STACK-V2 (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, the Ruff website hosts rule-specific documentation, including descriptions and examples, which we scrape to construct meta-task prompts $M_{I_{\mathcal{L}}}$ for each idiom. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix A.1.1.
3.2 Instruction Supervised Fine-Tuning
As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of 50 linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model’s output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix A.1.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some idioms, the final distribution is approximately 70:30 in favor of files with no violations. This results in a total of 53k training instances spanning 50 idioms.
3.3 Verifiable Reward Model and Preference Optimization
For preference optimization, we adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model. It samples $k$ outputs per input, computes rewards for them, and constructs contrastive win–loss pairs based on the reward distribution and a threshold $\eta$ (Figure 2). We detail the verifiable linter-based reward model and contrastive pair sampling procedure below.
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter’s line numbers (yellow circle in “Verifiable Reward Model”, Figure 1) as ground truth and the model’s predicted lines (blue circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$.
Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
$\Phi^{RL}=\arg\max\sum_{(x,y_{win},y_{lose})\in\mathcal{D}_{p}}\log\sigma\left(\beta\log\frac{\Phi^{RL}(y_{win}|x)}{\Phi^{SFT}(y_{win}|x)}-\beta\log\frac{\Phi^{RL}(y_{lose}|x)}{\Phi^{SFT}(y_{lose}|x)}\right)$
Here, $\sigma$ denotes the sigmoid function, and $\beta=0.1$ is the KL penalty coefficient, corresponding to low-to-moderate regularization.
3.4 Training with Reasoning Traces
Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning. To obtain CoT traces that help the LLM arrive at the correct answer, we adopt a rejection sampling approach (“Rejection Sampling SFT” in Figure 2) for collecting SFT data. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained LLM capable of CoT reasoning (e.g., Qwen3-4B), and compute the reward $r_{y_{i}}$ for each response, following a similar procedure as in RS-DPO. However, instead of forming contrastive pairs, we discard any $y_{i}$ with reward $r_{y_{i}}<\gamma$, where $\gamma=1$. In other words, we discard CoT–response pairs that are incorrect or improperly formatted and retain only those that are fully correct. The reward is applied only to the final response, obtained after extracting or parsing the CoT trace. We also remove any cases where the CoT fails to terminate or produce a final answer.
For some inputs $x$, no valid $y_{i}$ may be found. In such cases, we skip the input entirely. To promote meta-task diversity, we retain at most two valid responses $y_{i}$ per input. We sample multiple $y_{i}$ only for inputs with at least one violation, and only a single $y_{i}$ for inputs with no violations, this helps maintain a similar no-violation-to-violation ratio of 71:29 as in the regular SFT data. The latter case is also more likely to yield a valid CoT–response pair within the token limit. When more valid responses are available than needed, we keep the shortest $y_{i}$ completions, as they typically reflect more concise reasoning (the final answer usually consumes a similar number of tokens across samples). Following this policy, we collect 52.7k instances, which we use to train the base model using SFT.
After training the LLM on this dataset, we obtain a CoT-capable SFT model $\Phi_{CoT}^{SFT}$. We then apply the same RS-DPO procedure described in Section 3.3, with the only modification being that each $y_{i}$ now contains both the CoT trace and the final response.
4 Experiments
4.1 PEP Hard Idiom Benchmark
To evaluate whether MetaLint training helps the base LLM learn to interpret high-level idiom specifications, and apply this skill to harder, more abstract idioms that linters struggle to detect, we construct a benchmark of “hard idioms” sourced from PEPs. We begin by identifying fifteen PEPs that define semantic or abstract behaviors beyond what syntax or token patterns capture, then design specific heuristics for each to detect code that violates their guidelines. Using these heuristics, we search the STACK‑V2 corpus for violations, prioritizing high recall over strict precision to retrieve a broad set of candidates for manual selection. These “hard idioms” cannot be reliably detected by simple pattern matching, making them ideal for testing whether models truly understand idiom specifications or memorize frequent patterns seen in the Ruff idioms. From the retrieved examples, we handpick 15–20 representative files per PEP that clearly illustrate each violation. For each file, we identify the exact lines showing the violation (“before” snippet) and author a corrected version (“after” snippet) that complies with the PEP, forming the ground truth. To balance the benchmark, we include negative examples (code without violations) by selecting files retrieved for one PEP and confirming they don’t violate others. This results in 536 challenging examples, with 52% containing at least one violation and 48% being violation-free, allowing us to evaluate MetaLint’s ability to generalize from easy to hard idioms.
4.2 Evaluation Metrics
We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs. For both tasks, we compute precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each idiom, treating each idiom as a distinct class. Since the distribution of idiom occurrences may be imbalanced, we use macro averaging across idioms to weigh each class equally.
Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations. For each instance $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$, we define the indicator variable:
$\mathds{1}[x]=\begin{cases}1&\text{if }x\text{ is true}\\ 0&\text{otherwise}\end{cases}$
The idiom detection metrics per idiom $I$ are defined as:
$P_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|=0]\cdot\mathds{1}[|\hat{y}|>0]\right)}$
$R_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|=0]\right)}$
We compute the final detection metrics by macro averaging across all idioms:
$P_{\text{Det}}=\frac{1}{|I|}\sum_{I}P_{I},\quad R_{\text{Det}}=\frac{1}{|I|}\sum_{I}R_{I},\quad F_{\text{Det}}=\frac{2P_{\text{Det}}R_{\text{Det}}}{P_{\text{Det}}+R_{\text{Det}}}$
For idiom localization, we compute set level precision, recall, and F1-score for each instance $(x,y)\in\mathcal{D}$, comparing the predicted and ground truth violation lines:
$P_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|\hat{y}|},\quad R_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|y|},\quad F_{\text{Loc}}=\frac{2P_{\text{Loc}}R_{\text{Loc}}}{P_{\text{Loc}}+R_{\text{Loc}}}$
4.3 Baselines
We evaluate a range of state-of-the-art open and closed-source LLMs, trained for coding and reasoning tasks, on the PEP hard idiom benchmark to contextualize the improvements achieved by the MetaLint framework in flagging idiom violations from high-level specifications. Among open source models, we include instruction-tuned variants from the Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), and Qwen3 (Team, 2025) families. For closed-source models, we evaluate several OpenAI systems, including GPT-4o (Hurst et al., 2024), o3 mini, and o4 mini (OpenAI, 2025), and GPT-4.1 (OpenAI, 2025). We choose these models for their strong performance on a variety of code and reasoning benchmarks, and we evaluate whether code-specific training leads to measurable improvements. We also compare different model sizes, ranging from 3B and 4B to 32B parameters, in the open-source group
4.4 Evaluating Transfer Performance
To evaluate the extent to which the MetaLint training framework encourages adaptive reasoning over idiom specifications, as opposed to memorization, we construct a synthetic test set of 5.3k instances spanning 50 Ruff idioms, using the same data generation procedure described in Section 3.1. The data has a similar distribution of no-violation to violation examples as the SFT training set, with a 74:26 split in favor of no violations. We carefully select idioms to vary in their degree of overlap with those used for training, as shown in Figure 3. The selected idioms fall into three categories:
In-Domain: We select 5 idioms that are identical to those seen during SFT training, to assess whether MetaLint improves performance on idioms it was explicitly trained on.
Near Transfer: We include 10 idioms with specifications that closely resemble those in the training set but differ slightly. This category is crucial for probing memorization: if MetaLint relies too heavily on memorized training idioms, it may misclassify or underperform on these similar-looking test idioms due to interference.
Far Transfer: We choose 35 idioms that are substantially different from the training idioms to assess whether the LLM can rely solely on the provided specification. Success in this category would indicate that MetaLint fosters generalizable reasoning skills rather than dependence on specific training instances.
4.5 Evaluating Easy-to-Hard Generalization
To evaluate easy-to-hard generalization, we use the PEP hard idiom benchmark and compare the MetaLint-trained model against the baselines described in Section 4.3. The PEP idioms are held out for all models, including MetaLint, which is trained only on synthetic data generated from Ruff idioms. If training on the easier Ruff idioms leads to improved performance of MetaLint over the base models on the harder PEP idioms, it provides evidence that MetaLint encourages adaptive reasoning rather than memorization, thereby supporting easy-to-hard generalization.
5 Results
To test whether MetaLint enables the base LLM to interpret idiom specifications rather than simply memorize training examples, we evaluate its ability to generalize beyond the 50 Ruff idioms seen during training (Section 4.4) to both similarnear-transfer and significantly different far-transfer idioms. We also hypothesize that the RS-DPO (preference optimization) stage improves generalization over SFT alone, and that chain-of-thought (CoT) variants outperform their standard counterparts. Additionally, we assess easy-to-hard generalization by training on synthetic data for simpler Ruff idioms and evaluating on harder PEP idioms; if the model has truly learned to interpret high-level specifications, it should improve on these harder cases as well. Once again, we expect RS-DPO and CoT variants to yield stronger generalization. We present and analyze the results of these experiments in the following sections.
5.1 Ruff Idioms Performance
Table 1 presents Qwen3-4B’s idiom detection and localization performance on the full Ruff idiom transfer test set (Section 4.4) after MetaLint training, comparing CoT and non-CoT variants across both SFT and DPO stages. The untrained Qwen3-4B without CoT achieves decent detection precision but performs poorly on detection recall and localization. After SFT training, it shows substantial improvements in precision and all localization metrics, with modest recall gains. RS-DPO training further boosts detection recall and localization significantly (over 3× and nearly 5×, respectively), although with a slight drop in detection precision.
This precision drop stems from contrastive training examples where input $x$ has no violations, reducing the model’s accuracy in identifying such cases. We observe a strong correlation between the fraction of no-violation samples in training and detection precision. Including more such samples can improve precision but at the cost of recall and localization, leading to an unfavorable tradeoff. We find that including just 0.05% of no-violation data yields the best balance, minimizing precision loss while maximizing recall and improving detection F-score. Detailed ablation results are reported in Appendix A.2.2.
For the CoT variant, the untrained model already achieves strong detection precision and recall, along with high localization metrics—outperforming the SFT-only non-CoT model and trailing only slightly behind the SFT+RS-DPO non-CoT model (except in precision). To train the CoT model, we first apply RS-SFT to promote correct CoT reasoning and response formatting. This step is critical, as direct RS-DPO training alone causes metric collapse due to lack of absolute correctness supervision (since RS-DPO optimizes relative preference). RS-SFT leads to similar gains as non-CoT SFT, improving precision, recall, and localization. Final RS-SFT+RS-DPO training with 5% no-violation data mirrors the trends seen in the non-CoT case: increased recall, slight drop in precision, and overall gains in detection F-score and localization. Overall, the CoT model (RS-SFT+RS-DPO) is the strongest, achieving the highest precision and localization performance. However, the non-CoT SFT+RS-DPO model closely matches it in recall and approaches it on localization. Given that CoT inference is nearly 6× slower111measured on vLLM with 1× A100 80GB GPU the non-CoT model may be preferable in settings where detection recall and localization are the primary concerns.
5.2 Ruff Idioms Transfer Learning Performance
Table 2 presents MetaLint’s idiom detection performance across training setups for each transfer condition (in-domain, near-transfer and far-transfer) on the Ruff idiom benchmark. Untrained models perform poorly on in-domain idioms, especially in recall, suggesting these idioms are more subtle and difficult. Their stronger baseline performance on near and far transfer tasks likely reflects simpler or more surface-level idioms in those settings. SFT improves both precision and recall in-domain, but only precision improves in transfer settings, indicating limited generalization. Adding RS-DPO leads to consistent gains across all conditions, with particularly strong improvements in near and far transfer. This suggests that preference tuning encourages models to prioritize patterns that generalize beyond static rules.
CoT Untrained models already show high precision, likely due to the structured reasoning helping with correct interpretation of the idiom specifications and the output formatting instructions. RS-SFT improves recall with minimal changes to precision, though a small recall drop in near transfer may be due to a high baseline level of performance. Adding RS-DPO further boosts recall in all settings while maintaining strong precision. The largest recall improvements are seen in the in-domain setting, where idioms are hardest. These results offer some support for our claims that instruction-tuned models trained on synthetic linter signals generalize to novel idioms (the near and far transfer settings).
5.3 PEP Hard Idiom Benchmark Performance
Table 3 presents the results of evaluating the MetaLint DPO models and all baselines on the PEP hard idiom benchmark (a more comprehensive set of results is shown in Table 7). The MetaLint Qwen3-4B models significantly outperform the base Qwen3-4B models across all metrics in both the CoT and non-CoT settings, for both idiom violation detection and localization.
Interestingly, the non-CoT model achieves significantly higher detection recall and slightly higher F-score than the CoT variant, despite being less precise. Our analysis reveals that the CoT model’s lower detection recall results from its more conservative interpretation of the idiom specification and errors such as incorrect idiom interpretation, overthinking, and skipping lines, as detailed in Appendix A.3.3. In terms of detection F-score, the MetaLint models perform competitively, with only the closed-source GPT models, Qwen3-32B with CoT, and DeepSeek-R1-Distill-Qwen-32B with CoT surpassing the non-CoT model. Notably, the non-CoT model achieves the highest detection recall among all models evaluated, while the CoT model ranks among the best in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
For localization, the MetaLint models trail behind larger 32B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix A.3.2). This is notable given that the MetaLint models are much smaller (4B parameters) and trained only on synthetic data derived from easy idioms.
Overall, the strong performance, especially the best-in-class recall of the non-CoT model, demonstrates the effectiveness of our framework in achieving easy-to-hard generalization. This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction tuning, rather than relying on rote memorization.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Untrained 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
SFT+RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
CoT Untrained 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
CoT (RS-SFT) 0.935 0.8183 0.8727 0.6639 0.65 0.6569
CoT (RS-SFT + RS-DPO) 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Table 1: MetaLint Qwen3-4B results across training conditions evaluated on the Ruff meta-linting task.
Model In-Domain Near Transfer Far Transfer
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Untrained 0.45 0.14 0.22 0.58 0.24 0.34 0.54 0.29 0.38
SFT 0.93 (+0.48) 0.74 (+0.6) 0.83 (+0.61) 0.89 (+0.31) 0.24 (+0) 0.38 (+0.04) 0.72 (+0.18) 0.27 (-0.02) 0.39 (+0.01)
SFT+RS-DPO 0.72 (+0.27) 1 (+0.86) 0.83 (+0.61) 0.76 (+0.18) 0.8 (+0.56) 0.78 (+0.44) 0.75 (+0.21) 0.81 (+0.52) 0.78 (+0.4)
CoT Untrained 0.87 0.5 0.63 0.95 0.88 0.91 0.87 0.68 0.76
CoT (RS-SFT) 0.87 (+0) 0.73 (+0.23) 0.8 (+0.17) 0.97 (+0.02) 0.86 (-0.02) 0.91 (+0) 0.94 (+0.07) 0.82 (+0.14) 0.88 (+0.12)
CoT (RS-SFT + RS-DPO) 0.86 (-0.1) 0.85 (+0.35) 0.85 (+0.22) 0.97 (+0.02) 0.92 (+0.04) 0.94 (+0.03) 0.92 (+0.05) 0.86 (+0.18) 0.89 (+0.13)
Table 2: MetaLint Qwen3-4B detection results across training conditions evaluated on the Ruff meta-linting task across transfer conditions (In-Domain, Near Transfer and Far Transfer)
Detection Localization
Model $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B with CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
Qwen3-14B 0.9021 0.4612 0.6103 0.289 0.2521 0.2693
Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
Qwen3-4B MetaLint (SFT+RS-DPO) (Ours) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) (Ours) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
Table 3: Results on the hard PEP benchmark to measure easy to hard generalization.
6 Conclusion
Our results demonstrate that MetaLint may encourage more adaptive reasoning over idiom specifications instead of rote memorization of code patterns through instruction-following and easy-to-hard generalization from synthetic data. By training on synthetic examples of easy-to-detect idioms derived from linter signals, MetaLint enables both Qwen3-4B CoT and non-CoT variants to generalize well across Ruff idioms in in-domain, near-transfer, and far-transfer settings, as well as to real-world, hard-to-detect PEP idioms. The models show substantial improvements in both detection and localization, particularly after RS-DPO training, with the non-CoT variant achieving the highest detection recall among all evaluated models, despite its smaller size and training solely on synthetic data. Additionally, it achieves this with just 4B parameters and no test-time compute in the non-CoT setting, making it one of the most efficient models. These findings underscore the effectiveness of synthetic instruction-tuned training for fostering generalization and reasoning, even in low-resource scenarios where annotated examples of harder idiom categories are limited or unavailable. Finally, we note that for easy idioms that are more “mechanical” to detect, linters remain the most cost-effective option. However, LLMs—when trained with the MetaLint approach—enable detection of more abstract idioms, potentially unlocking support for personalized linting, and offering a no-code solution for code quality enforcement.
7 Limitations and Future Work
Despite the promising results achieved by MetaLint, our work has several limitations that we plan to address in future research. First, our experiments are restricted to a single model family (Qwen3) and model size (4B). We intend to extend our evaluations to other model families, including LLaMA and Phi, as well as larger Qwen models (8B and 14B). In the CoT setting, it remains an open question whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model. We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data. As a result, we adopted a simpler rejection sampling strategy. Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as direct GRPO (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data. Our current experiments also focus exclusively on the Python programming language. Future work will explore extensions to other languages such as Java using Java-specific linters, as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa. Finally, while we prompt and train the models to generate refactoring fixes for code with idiom violations to improve code quality, we do not evaluate this capability in the present work and leave it for future exploration.
Author Contributions
Atharva Naik: Conception, Ruff Idiom Data Generation, SFT, RS-SFT, RS-DPO Experiments, Baselines, Writing
Lawanya Baghel: RS-DPO Experiments, PEP Hard Idiom Benchmark Creation, Writing
Dhakshin Govindarajan: PEP Hard Idiom Benchmark Creation
Darsh Agrawal: PEP Hard Idiom Benchmark Creation
Daniel Fried: Advising, Conception, Writing
Carolyn Rose: Advising, Conception, Writing
Acknowledgments
This work was partially supported by a generous collaborative research grant from Oracle Labs. We are grateful to Jack Sullivan, Qinlan Shen, Adam Pocock, and Nidhi Chandra at Oracle Labs for their guidance, feedback, and insightful discussions that played a key role in shaping the core ideas behind MetaLint and advancing the project.
Appendix A Appendix
A.1 Method Additional Details
A.1.1 MetaLint Instruction Following Prompt
We used the following instruction following style prompt to train the model with synthetic Ruff idiom data for the meta-linting task:
MetaLint Instruction Following Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} Violations per idiom:
An example input with the code file and idiom spec populated as well as the expected JSON style output is shown below:
Example Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations. Rationale: Type annotations are a good way to document the return types of functions. They also help catch bugs, when used alongside a type checker, by ensuring that the types of any returned values, and the types expected by callers, match expectation. Example:
def _add(a, b):
    return a + b
Use instead:
def _add(a: int, b: int) -> int:
    return a + b
Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say ’NO VIOLATIONS FOUND’ if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file:
  1 # -*- coding: utf-8 -*-
  2 # pragma pylint: disable=unused-argument, no-self-use
 86     def _reload(self, event, opts):
 87         """Configuration options have changed,
    save new values"""
 88         self.options = opts.get("fn_cisco_amp4ep", {})
 89         validate_opts(self)
 90
 91     @function("fn_amp_move_computer")
 92     def _fn_amp_move_computer_function(self, event, *args,
    **kwargs):
 93         """Function: Move computer to a group with given
    connector guid and group guid."""
 94         try:
Violations per idiom:
Example Output **Idiom ANN202 Violations:**
{"line": " 86     def _reload(self, event, opts):", "fix": null}
{"line": " 92     def _fn_amp_move_computer_function(self,
event, *args, **kwargs):", "fix": null}
A.1.2 DPO Contrastive Pair and RS-SFT Sampling Details
To generate RS-DPO contrastive samples (or RS-SFT outputs) from the baseline SFT (or untrained) models, we used the following hyperparameters: nucleus sampling with a maximum of 2048 new tokens, $k=5$ sampled outputs per input, temperatures picked cyclically from {0, 0.3, 0.5, 0.7, 1}, a top-$p$ (cumulative probability threshold) of 0.95, and a seed of $42+i$, where $i\in\{1,\dots,k\}$, to encourage both reproducibility and output diversity.
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded “Baseline Inference Prompt” described in Section A.1.4.
A.1.3 Training Hyperparameters and Computational Environment
We use the following hyperparameters for the SFT and DPO training experiments.
SFT/RS-SFT hyperparameters:
For both SFT and RS-SFT we fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision. The model is trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and a warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled to reduce memory usage, with non-reentrant mode. Evaluation is performed every 2000 steps, and checkpoints are saved at the same interval. Special tokens are manually handled in the chat template without automatic insertion. The training uses 12 preprocessing workers and is seeded with 42 for reproducibility.
RS-DPO parameters:
We fine-tune the model using RS-DPO with bfloat16 precision and a reward shaping parameter $\beta=0.1$. Training is performed for 1 epoch with a learning rate of 5e-7, cosine learning rate scheduling, and a warmup ratio of 0.1. We use a maximum input length of 3500 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled with non-reentrant mode to optimize memory usage. The optimizer is AdamW, and evaluation is conducted every 200 steps with checkpoints saved at the same interval. The training is seeded with 42 for reproducibility.
Computational Environment: All SFT, RS-SFT and RS-DPO experiments were conducted on a Linux server equipped with two NVIDIA A100 80GB PCIe GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03. Each job had access to 100 GB of CPU memory and 2 CPU cores. Training was performed using mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage. For inference we used a similar environment but varied number of GPUs based on model sizes and needs.
A.1.4 Baseline Inference Details
We use the following hyperparameters for performing inference with the baseline LLMs:
Open Source LLMs: We perform nucleus sampling with 8192 max-new tokens, temperature of 0.7, top-p (cumulative probability threshold) of 0.95 and seed of 42 (to promote reproducibility).
Closed Source LLMs: We use the chat completion OpenAI API with max tokens of 1024 for GPT-4.1 and GPT-4o and max completion tokens of 3000 for o3-mini and o4-mini. We use default parameters for everything else (temperature of 1 and top-p of 1, no presence penalty).
Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches. For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Idiom Benchmark.
Baseline Inference Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called “### Final Idiom Violations Found”. Structure you response for a given idiom XYZ as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:**
[…]
Inquiry.objects.filter(owner=kwargs.pop(’user’)))", "fix": null}
[…]
4 from inquiry.models import *",
[…]
and as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:** NO VIOLATIONS FOUND Violations per idiom:
A.2 Additional Experimental Details
A.2.1 Idioms Chosen for Ruff Idiom Transfer Dataset
Table 4 lists the Ruff idioms used in the SFT training and synthetic transfer evaluation test sets. Idioms are grouped by their source linter and cover a range of syntax, semantics, naming, and upgrade-related rules.
Training Set Idioms Test Set Idioms
PyFlakes: PyFlakes:
F405, F501, F502, F601, F621 F403, F406, F503, F602, F622
pycodestyle: pycodestyle:
E402, E701, E721, E741, E743 E401, E702, E722, E731, E742
Naming: Miscellaneous:
N801, N802, N803, N804, N805, ERA001, C901, I001, I002, BLE001
N806, N807, N811, N812, N813 (shared with training)
pyupgrade: flake8 annotations:
UP001, UP002, UP003, UP004, UP005, UP006, ANN001, ANN002, ANN003, ANN201, ANN202,
UP007, UP008, UP009, UP010, UP011, ANN204, ANN205, ANN206
UP040, UP044, UP045, UP046, UP047
Miscellaneous: flake8 async:
ERA001, C901, I001, I002, BLE001 ASYNC100, ASYNC105, ASYNC109, ASYNC110,
ASYNC115, ASYNC116, ASYNC210, ASYNC220,
ASYNC221, ASYNC222, ASYNC230, ASYNC251
Bugbear: flake8 bandit:
B002, B003, B004, B005, B006, S102, S103, S104, S105, S106,
B007, B008, B009, B010, B012 S107, S108, S110, S112, S113,
S201, S202, S301, S302, S303
Table 4: Ruff idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
Figure 3: ID: In-Domain, NeT: Near Transfer, FaT: Far Transfer.
A.2.2 DPO No Violation Fraction Ablations
We analyze the impact of varying the amount of samples with zero violations used for RS-DPO training. These experiments were motivated by initial findings comparing models trained only on data with at least one violation to those trained on the full dataset. By design, RS-DPO generates significantly more training data for cases with at least one violation, due to greater variance in reward signals. This is further amplified by the fact that the initial SFT policy/checkpoint is already quite accurate in handling cases with NO VIOLATIONS FOUND leading to low variance in reward across responses.
Our early experiments showed that excluding all NO VIOLATIONS FOUND cases led to notable gains in recall and line-level localization. However, this came at the cost of a significant drop in precision compared to the SFT policy/base model. Further analysis revealed a sharp decline in the accuracy of predicting NO VIOLATIONS FOUND, from nearly 99% down to 70-80%, with performance worsening monotonically over training steps. Conversely, training on the full dataset (i.e., including 100% of the NO VIOLATIONS FOUND cases) improved precision but offered only modest gains in recall and localization, which also degraded with continued training. These findings suggest that while some NO VIOLATIONS FOUND data is necessary to maintain high precision, too much of it may hinder recall and localization.
To investigate this trade-off, we experimented with keeping only a fraction of the NO VIOLATIONS FOUND data during training. Specifically, we randomly sampled $k$of such data, varying $k$ across {0%, 2%, 5%, 10%, 20%, 40%, 100%}. These percentages were selected based on observed trends: 20%, 40%, and 100% yielded similar results, which discouraged further tests at 60% or 80%, while 2% and 5% were chosen due to a noticeable performance jump between 0% and 10%. We found that 5% offered a favorable middle ground, largely retaining or slightly reducing precision, while preserving most of the recall (resulting in the highest detection F-score), and only modestly impacting line-level localization. Based on these insights, we conducted a limited ablation on the CoT model, evaluating 2% and 5% inclusion to determine the optimal setting for both detection and localization (as shown in Table 6).
Fraction of NV data Det. Line overlap (Det. + Loc.)
P R F P R F
0% 0.6268 0.9577 0.7577 0.6777 0.6932 0.6854
2% 0.671 0.9128 0.7734 0.6681 0.6812 0.6746
5% 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
10% 0.7584 0.8114 0.784 0.6263 0.6474 0.6367
20% 0.8382 0.7227 0.7762 0.5721 0.5815 0.5768
40% 0.8683 0.5618 0.6822 0.4683 0.4735 0.4709
100% 0.8565 0.4152 0.5593 0.4041 0.4056 0.4048
Table 5: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT. Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND. Conversely, including 100% improves precision but leads to reduced recall and localization performance. All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Idiom Transfer test set.
Fraction of NV data Det. Line overlap (Det. + Loc.)
P R F P R F
2% 0.9226 0.8901 0.906 0.7688 0.7638 0.7663
5% 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Table 6: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model with CoT. We perform limited ablations because of the insights from the non CoT model training.
A.3 More Results
A.3.1 Expanded Results on the PEP Hard Idiom Benchmark
We show the expanded results across various model sizes for the evaluated model families in Table 7. We note that most results follow the expected trends with more parameters or CoT usage leading to better performance but there are soem exceptions to the trend. We mainly see this for cases like Qwen2.5 and Qwen2.5Coder families. We note that Qwen2.5Coder-7B-Instruct has almost zero metrics because it always predicts NO VIOLATIONS FOUND for all instances and Qwen2.5Coder-14B-Instruct has really low scores because of similar reasons. for Qwen2.5 family we notice that 32B variant performs a bit worse than 32B.
We also analyze MetaLint SFT models on the hard PEP benchmark and observe that they perform similarly or slightly worse than the base untrained models. This suggests that SFT alone may lead to overfitting on the Ruff idiom distribution and struggles to generalize from easy to hard cases without DPO training. These findings highlight the importance of the DPO (preference-tuning) stage in the MetaLint pipeline. However, we also emphasize that while the SFT stage can limit generalization, it remains essential for effective DPO training, as it teaches the LLM to follow the correct output format and establishes a strong base policy. This is supported by our experiments with the CoT model, where applying RS-DPO directly to the Qwen/Qwen3-4B model (without SFT) led to near-zero performance across all metrics, as the model consistently failed to produce outputs in the required format.
Detection Localization
Model P R F P R F
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B with CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
Qwen3-14B 0.9021 0.4612 0.6103 0.289 0.2521 0.2693
Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
Qwen2.5-3B-Instruct 0.0667 0.0033 0.0063 0.0036 0.0036 0.0036
Qwen2.5-7B-Instruct 0.4333 0.1379 0.2092 0.0585 0.0518 0.0549
Qwen2.5-14B-Instruct 0.8017 0.4324 0.5618 0.2389 0.2158 0.2267
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
Qwen2.5Coder-3B-Instruct 0.7802 0.411 0.5384 0.1257 0.0745 0.0936
Qwen2.5Coder-7B-Instruct 0.0667 0.0033 0.0063 0 0 0
Qwen2.5Coder-14B-Instruct 0.2 0.0443 0.0726 0.0294 0.0264 0.0278
Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
DeepSeek-R1-Distill-Qwen-7B with CoT 0.7143 0.2841 0.4065 0.1064 0.1122 0.1092
DeepSeek-R1-Distill-Qwen-14B with CoT 0.69 0.2345 0.35 0.1856 0.1245 0.149
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
MetaLint (SFT) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
MetaLint w CoT (RS-SFT) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
Table 7: Result on the hard PEP benchmark to measure easy to hard generalization.
A.3.2 Statistical Significance of Results on the PEP Hard Idiom Benchmark
To analyze the statistical significance of performance differences over the PEP benchmark, we conduct Wilcoxon signed-rank tests comparing various MetaLint variants against each other and against baseline models. We evaluate instance-level detection accuracy (binary labels indicating whether the LLM correctly predicted the presence of a violation) as well as instance-level precision and recall for line-level localization. To control for multiple comparisons, we apply a Bonferroni correction to adjust the significance threshold $\alpha$ as $\alpha=\frac{0.05}{m}$ where $m$ is the number of comparisons (or rows in any given statistical significance table in this case).
Table 9 reports the Wilcoxon signed-rank test statistic and corresponding $p$-value (in parentheses) for detection accuracy, localization precision, and localization recall when comparing various MetaLint variants to assess the effects of RS-DPO and CoT. We find that applying RS-DPO to the base SFT policy leads to statistically significant improvements in both detection and localization performance, with RS-DPO consistently outperforming the original SFT checkpoint across all three metrics with it being always better for localization. For the CoT variant, RS-DPO also yields consistent but less significant gains, likely because the RS-SFT CoT checkpoint is already relatively strong. Finally, we observe no statistically significant difference between the CoT (RS-SFT+RS-DPO) and the standard (SFT+RS-DPO) variant, suggesting that CoT does not provide a meaningful additional benefit in this setting.
Table 10 shows the statistical significance of comparing the base untrained model Qwen3-4B with its MetaLint variants (SFT and SFT+RS-DPO), and the Qwen3-4B CoT model with MetaLint w/ CoT (RS-SFT and RS-SFT+RS-DPO). The SFT variant yields significant gains in detection and localization recall, but not in localization precision. The SFT+RS-DPO model improves significantly across all three metrics. In contrast, training RS-SFT from the Qwen3-4B w/ CoT base does not yield significant improvements. However, the RS-SFT+RS-DPO variant produces significant gains in localization precision and recall, but not detection. These results suggest that while SFT alone offers limited generalization, combining it with DPO reliably improves localization and can significantly boost detection when starting from a weaker base model.
Table 11 shows the statistical significance results when comparing the MetaLint (SFT+RS-DPO) and MetaLint w CoT (RS-SFT+RS-DPO) variants against various baselines. Here we want to higlight that MetaLint offers comparable performance across two out of three or all three metrics against several 32B models that outperform it like Qwen3-32B, Qwen3-32B w CoT, Qwen2.5Coder-32B and R1-Distill-Qwen-32B. Also the MetaLint non CoT (SFT+RS-DPO) variant has no singificant difference in performance compared to o3-mini, soldifying that MetaLint without CoT has generalized to the point of being as capable as o3-mini (even though the Qwen3-4B models without CoT and Qwen3-4B model with CoT perform worse than it with the difference being statistically singificant in Table 8).
Model Comparison Detection Localization P Localization R
Qwen3-4B vs o3-mini 1266.5 (7.20e-21) 743.5 (2.96e-11) 739.0 (4.23e-09)
Qwen3-4B w CoT vs o3-mini 921.5 (3.23e-09) 2385.5 (9.61e-02) 1891.0 (4.99e-04)
Table 8: Wilcoxon signed-rank test results comparing untrained Qwen3-4B variants with o3-mini, using Bonferroni-adjusted significance threshold $\alpha=0.025$. Each cell reports the test statistic (p-value).
Table 12 shows the effect of using a CoT for the Qwen3 model families and we notice that using a CoT leads to singificant gains for all metrics for the 4B and 8B models indicating that for smaller models CoTs might be essential for good performance on this task. However the 14B and 32B model only show statistically significant improvement in localization precision with the CoT indicating that the CoT might offer limited benefit for larger models.
Table 13 shows the effect of varying model scale for the Qwen3, Qwen2.5, Qwen2.5Coder, and DeepSeek-R1-Distill-Qwen families. For Qwen3 we see benefits moving from 4B to 8B abd 8B to 14B but no statistically significant difference moving from 14B to 32B when not using a CoT. Wehn using a CoT for Qwen3 we notice that the performance differences are rarely different in terms of statistical significant except for localizaiton performance between 4B and 8B and 8B and 14B. For R1-Distill-Qwen family we notice a significant difference moving from 14B to 32B but not for 7B to 14B. For the Qwen2.5Coder family we notice difference across all model scales, but the trend is weird with a big drop in performance from 3B to 7B and then a slow climb back to great performance around 32B. We notice that for the Qwen2.5 family which shows relatively reasonable trends with model scale, the performance differences are statistically singificant execpt for the performance gain from 14B to 32B being significant only for recall. To conclude the trends across model scales vary a lot across model families but in general the model size does help but differences may be smaller if the models are capable of reasoning and use a CoT.
Table 14 shows comparison between the GPT models. We only compared GPT-4o and its successor GPT-4.1 and o3-mini against o4-mini and the results show that GPT-4.1 is only significantly better for localization recall while o4-mini is beter than o3-mini for overall localization but not for detection.
Model Comparison Detection Localization P Localization R
MetaLint (SFT) vs MetaLint (SFT+RS-DPO) 7192.0 (1.92e-12) 0.0 (2.49e-20) 0.0 (2.92e-18)
MetaLint w CoT (RS-SFT) vs MetaLint w CoT (RS-SFT+RS-DPO) 839.5 (2.18e-03) 740.0 (3.95e-03) 523.0 (2.34e-05)
MetaLint (SFT) vs MetaLint w CoT (RS-SFT) 528.0 (6.91e-14) 11.0 (1.38e-15) 113.0 (7.95e-12)
MetaLint (SFT+RS-DPO) vs MetaLint w CoT (RS-SFT+RS-DPO) 8140.0 (5.55e-01) 2568.5 (8.42e-01) 2544.0 (4.44e-01)
Table 9: Wilcoxon signed-rank test results comparing MetaLint variants. Each cell reports test statistic (p-value). All the MetaLint models are trained Qwen3-4B variants. We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-4B MetaLint (SFT) 560.5 (8.64e-03) 363.5 (1.82e-02) 238.0 (4.85e-04)
Qwen3-4B vs Qwen3-4B MetaLint (SFT+RS-DPO) 7260.0 (4.13e-09) 411.0 (1.99e-15) 979.5 (7.30e-09)
Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT) 1224.0 (7.22e-01) 937.0 (6.12e-01) 918.5 (5.39e-01)
Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT+RS-DPO) 1728.0 (1.83e-02) 1011.0 (1.53e-03) 966.0 (1.02e-03)
Table 10: Wilcoxon signed-rank test results comparing MetaLint models against their untrained counterparts, with Bonferroni-adjusted significance threshold $\alpha=0.0125$. Each cell reports the test statistic (p-value).
Model Comparison Detection Localization P Localization R
Qwen3-8B vs MetaLint (SFT+RS-DPO) 8140.5 (7.21e-03) 1309.5 (2.19e-08) 2067.0 (2.12e-03)
Qwen3-8B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2070.0 (9.17e-01) 1974.5 (1.88e-01) 2161.0 (6.58e-01)
Qwen3-14B vs MetaLint (SFT+RS-DPO) 7304.0 (4.96e-01) 2816.5 (3.15e-02) 3159.0 (2.88e-02)
Qwen3-14B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2392.0 (2.78e-01) 2749.5 (1.26e-01) 2319.0 (1.32e-03)
Qwen3-32B vs MetaLint (SFT+RS-DPO) 7175.0 (4.48e-01) 3262.0 (1.93e-02) 2818.5 (5.56e-03)
Qwen3-32B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2677.5 (9.95e-03) 3479.0 (8.38e-02) 3180.5 (5.64e-04)
R1-Distill-Qwen-7B vs MetaLint (SFT+RS-DPO) 8244.0 (2.65e-08) 555.0 (8.12e-15) 1924.0 (1.91e-04)
R1-Distill-Qwen-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 2907.0 (7.07e-10) 915.5 (1.04e-12) 1569.5 (8.64e-06)
R1-Distill-Qwen-14B vs MetaLint (SFT+RS-DPO) 9877.0 (3.99e-06) 2582.5 (9.36e-06) 3085.0 (2.00e-03)
R1-Distill-Qwen-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 2660.0 (9.11e-08) 1703.0 (2.98e-06) 1791.0 (3.97e-05)
R1-Distill-Qwen-32B vs MetaLint (SFT+RS-DPO) 8677.5 (2.51e-01) 5767.5 (1.93e-01) 3705.0 (6.67e-06)
R1-Distill-Qwen-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 3125.0 (3.11e-02) 3641.5 (6.35e-02) 2175.0 (4.99e-06)
Qwen2.5-3B vs MetaLint (SFT+RS-DPO) 8001.0 (1.41e-15) 0.0 (1.24e-22) 68.5 (1.26e-19)
Qwen2.5-3B vs MetaLint w CoT (RS-SFT+RS-DPO) 949.0 (4.96e-23) 0.0 (4.24e-22) 0.0 (9.89e-20)
Qwen2.5-7B vs MetaLint (SFT+RS-DPO) 7312.5 (3.37e-10) 208.0 (1.44e-18) 610.0 (1.99e-12)
Qwen2.5-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 1187.5 (1.14e-14) 226.0 (2.60e-18) 406.5 (5.69e-14)
Qwen2.5-14B vs MetaLint (SFT+RS-DPO) 8677.5 (2.51e-01) 3045.5 (5.70e-04) 4006.0 (3.83e-01)
Qwen2.5-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 4123.0 (4.86e-01) 3228.0 (1.51e-03) 4383.0 (9.89e-01)
Qwen2.5-32B vs MetaLint (SFT+RS-DPO) 8640.0 (1.76e-04) 1492.5 (3.12e-08) 2971.5 (4.55e-02)
Qwen2.5-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 1792.0 (8.16e-06) 1166.0 (2.43e-08) 1983.5 (4.71e-03)
Qwen2.5Coder-3B vs MetaLint (SFT+RS-DPO) 11184.0 (8.64e-03) 953.5 (3.73e-12) 1716.0 (3.00e-07)
Qwen2.5Coder-3B vs MetaLint w CoT (RS-SFT+RS-DPO) 3683.5 (6.45e-03) 1126.0 (4.39e-11) 1403.5 (1.32e-08)
Qwen2.5Coder-7B vs MetaLint (SFT+RS-DPO) 8001.0 (1.41e-15) 0.0 (7.69e-23) 0.0 (2.02e-20)
Qwen2.5Coder-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 949.0 (4.96e-23) 0.0 (2.61e-22) 0.0 (6.55e-20)
Qwen2.5Coder-14B vs MetaLint (SFT+RS-DPO) 9123.5 (1.04e-12) 289.0 (8.19e-20) 736.0 (7.32e-14)
Qwen2.5Coder-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 1112.0 (1.82e-19) 159.5 (7.63e-20) 408.5 (3.02e-15)
Qwen2.5Coder-32B vs MetaLint (SFT+RS-DPO) 6833.5 (2.86e-01) 4500.0 (9.59e-01) 2651.5 (7.07e-05)
Qwen2.5Coder-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 1039.5 (1.16e-02) 2655.0 (9.39e-01) 1235.5 (1.44e-05)
o3-mini vs MetaLint (SFT+RS-DPO) 7520.0 (4.83e-02) 4986.0 (5.20e-01) 4427.5 (2.87e-01)
o3-mini vs MetaLint w CoT (RS-SFT+RS-DPO) 1944.0 (7.15e-04) 3169.0 (5.16e-01) 2683.0 (4.23e-01)
Table 11: Wilcoxon signed-rank test statistics and p-values comparing MetaLint variants against baseline models. All the MetaLint variants are Qwen3-4B variants and Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0017$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-4B w CoT 1260.0 (3.99e-08) 425.0 (1.91e-09) 624.5 (2.25e-05)
Qwen3-8B vs Qwen3-8B w CoT 1924.0 (4.27e-03) 1132.5 (8.69e-06) 1005.0 (1.15e-04)
Qwen3-14B vs Qwen3-14B w CoT 1691.0 (2.01e-01) 2127.0 (4.27e-04) 2398.5 (1.26e-01)
Qwen3-32B vs Qwen3-32B w CoT 1572.5 (2.75e-01) 1767.0 (6.35e-06) 2596.5 (7.31e-02)
Table 12: Wilcoxon signed-rank test results measuring the effect of Chain-of-Thought (CoT) prompting across Qwen3 model scales. Each cell reports test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-8B 546.0 (2.35e-08) 618.0 (1.37e-04) 572.5 (1.72e-03)
Qwen3-8B vs Qwen3-14B 1350.0 (2.11e-03) 1503.5 (3.96e-05) 1008.5 (2.69e-07)
Qwen3-14B vs Qwen3-32B 875.0 (2.22e-02) 3129.5 (7.94e-01) 2061.0 (4.14e-01)
Qwen3-4B w CoT vs Qwen3-8B w CoT 1468.5 (1.90e-02) 1578.5 (1.07e-01) 1081.5 (2.60e-03)
Qwen3-8B w CoT vs Qwen3-14B w CoT 904.5 (1.40e-01) 1248.0 (1.66e-03) 1099.5 (1.82e-05)
Qwen3-14B w CoT vs Qwen3-32B w CoT 850.0 (3.78e-02) 1834.5 (4.87e-01) 2396.0 (4.40e-01)
R1-Distill-Qwen-7B vs R1-Distill-Qwen-14B 4278.0 (2.67e-01) 1431.0 (4.66e-03) 1962.5 (6.01e-01)
R1-Distill-Qwen-14B vs R1-Distill-Qwen-32B 2432.0 (1.44e-12) 1475.5 (8.07e-11) 843.5 (6.38e-15)
Qwen2.5Coder-3B vs Qwen2.5Coder-7B 1541.0 (4.56e-14) 0.0 (3.46e-10) 0.0 (7.07e-10)
Qwen2.5Coder-7B vs Qwen2.5Coder-14B 8.0 (7.89e-04) 0.0 (2.04e-03) 0.0 (2.14e-03)
Qwen2.5Coder-14B vs Qwen2.5Coder-32B 423.0 (2.83e-27) 100.5 (2.00e-22) 43.5 (9.29e-23)
Qwen2.5-3B vs Qwen2.5-7B 18.0 (2.43e-08) 0.0 (3.43e-05) 0.0 (3.58e-05)
Qwen2.5-7B vs Qwen2.5-14B 960.5 (1.70e-13) 500.5 (1.65e-11) 574.0 (2.62e-10)
Qwen2.5-14B vs Qwen2.5-32B 1925.0 (1.87e-04) 1858.5 (6.60e-03) 2108.5 (1.43e-02)
Table 13: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings. Each cell shows the test statistic (p-value). All Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
Model Comparison Detection Localization P Localization R
GPT-4o vs GPT-4.1 2550.0 (9.21e-01) 3961.0 (5.83e-01) 2207.5 (4.48e-06)
o3-mini vs o4-mini 575.0 (6.68e-01) 1171.0 (1.44e-03) 1079.0 (1.47e-04)
Table 14: Wilcoxon signed-rank test results comparing GPT model variants. Each cell shows the test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.025$.
A.3.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
We observe that a significant portion of the lower detection recall of the CoT MetaLint Qwen3-4B model, relative to its non CoT counterpart, can be attributed to its higher tendency to predict NO VIOLATIONS FOUND in cases that do, in fact, contain violations. Specifically, the CoT model fails to flag violations in 89 additional instances compared to the non CoT model, amounting to nearly 17% of the evaluation set (89 out of 536 examples).
The idiom wise distribution of these missed violations is shown in Figure 4. While the failure distribution follows a somewhat long tail pattern, the most significant drops occur for PEP 614, PEP 616, and PEP 593. Notably, if the CoT model matched the non CoT model’s performance on just these three PEPs, its detection recall would rise to 0.605, surpassing that of all open source baselines evaluated.
Upon inspecting CoT traces for these and other idioms (see examples in Table 15), we identify several recurring failure modes: 1) Ambiguity in interpreting the idiom specification. For example, in PEP 614, which targets decorators with complex expressions, the CoT model often labels expressions that humans consider complex as simple. 2) Overthinking and repetitive reasoning traces, particularly for PEP 616. 3) Skipping or entirely missing lines that contain violations, again observed in PEP 616. 4) Underspecified idioms. For instance, in PEP 593, which recommends using the Annotated type from the typing module to attach metadata to type hints, the spec lacks clarity and concrete examples, making it hard to learn what constitutes a violation.
We also find similar issues in idioms like PEP 487, which discourages the use of metaclasses for simple customization tasks that could be handled via __init_subclass__ or __set_name__. The CoT model often misclassifies such “simple” use cases as complex.
Overall, these patterns suggest that the CoT model applies the idiom specifications more conservatively, resulting in higher precision but at the cost of reduced recall.
Figure 4: Distribution of comparative failures of the CoT MetaLint Qwen3-4B model relative to its non-CoT variant. While errors span a long tail across many PEPs, the majority are concentrated in three: PEP614, PEP593, and PEP616, which motivates our focused analysis on these cases.
PEP Chain of Thought Reason for Failure
614
… I need to check the code file for violations of the 614 idiom,
which is about using complex expressions in decorators. …
… The register variable is defined on line 7 as partial(register,
server=hippa). So, the decorator is a partial function. …
… the idiom allows for complex expressions in the decorator.
For example, if the decorator was something like
@lambda func: timed(…)(func), that would be allowed. …
… the register is a partial function, which is a simple expression. …
… the decorator is just a variable that’s a partial function. …
Ambiguity about what constitutes
a complex expression. The idiom
targets decorators with complex
expressions that make code
more verbose
525
… The user … wants me to check for … idiom, 525, which
is about using asynchronous generators with ”async for”
syntax … the violation would be if the code uses
synchronous generators (like regular for loops) instead of
asynchronous generators (using async for) … Wait, the
example given in the idiom is a class that uses __aiter__
and __anext__ to create an asynchronous iterator. The fix
is to use a function with async def and yield. So, I need to
check if there are any instances where the code is using
synchronous generators (like for loops) instead of
asynchronous ones. … Looking through the code, I
don’t see any for loops that are using ’async for’. …
CoT accidentally looks for the
incorrect construct at a crucial
point (async for instead of for)
487
… using __init_subclass__ and __set_name__ for class
customization without metaclasses … using metaclasses
for simple customization tasks makes code unnecessarily
complex … there’s a class called FirestoreClientMeta,
which is a metaclass … used for registration of transports
… the problem is whether this is a simple customization …
not just for a single customization but for managing
multiple transports … the idiom is about using metaclasses
for simple customization … but in this case … handling a
more complex task … maybe this is not a violation …
Ambiguity around what constitutes
a simple class. The idiom targets
instances where metaclasses
are used for simple customization
tasks that could have been handled
with ”__init_subclass__” or
”__set_name__”)
Table 15: Example chains of thought for various PEPs where the CoT model incorrectly flags NO VIOLATIONS FOUND instead of the non CoT model.