### v3

# MetaLint: Generalizable Idiomatic Code Quality Analysis through Instruction-Following and Easy-to-Hard Generalization

Abstract
Large Language Models excel at code generation but struggle with code quality analysis, where best practices evolve and cannot be fully captured by static training data. We introduce MetaLint, a training framework that treats code quality analysis as detecting best practice violations from high-level specifications over semantic code fragments (code idioms). Instead of training on a fixed set of rules, MetaLint reorganizes supervision around dynamically specified best practices using synthetic linter-derived labels, integrated with instruction-following and preference optimization. This encourages extrapolation to more complex, unseen best practices at test time, consistent with easy-to-hard generalization without retraining. To evaluate MetaLint, we create a new benchmark of hard-to-detect best practices inspired by Python Enhancement Proposals. Across this benchmark, MetaLint improves generalization to unseen best practices. Qwen3-4B achieves a 2.7× detection F-score gain (25.9% → 70.4%), the highest recall, and a 26.7% localization F-score, matching larger models such as o3-mini. These gains generalize across programming languages, model families, scales, reasoning settings, and linter sources.
Easy to Hard Generalization, Transfer Learning
1 Introduction
Code quality analysis offers a unique opportunity to study generalization in Large Language Models (LLMs) because best practices are contextual, evolving, and poorly captured by static training data. With the rise of LLMs for code, concerns around the quality of generated code, such as readability, maintainability, efficiency, and security, have become increasingly prominent (Singhal et al., 2024; Zheng et al., 2024). Researchers have explored LLMs’ potential to improve code quality through benchmarks, code review agents, and static analysis (Singhal et al., 2024; Vijayvergiya et al., 2024; Fang et al., 2025). However, studies show that LLMs often struggle with these tasks, and improvements via prompting or training are limited by static datasets reflecting narrow or outdated best practices (Vijayvergiya et al., 2024; Khare et al., 2023; Holden & Kahani, 2024; Zhang et al., 2024b). As a result, these systems often perform poorly on rare best practices, different code distributions (Holden & Kahani, 2024) or over-flag outdated best practices, leading to a negative user experience and wasted time (Vijayvergiya et al., 2024). This difficulty stems not only from limited or outdated training data, but also from supervision: models are trained on a fixed set of best-practice labels, which encourages memorization of individual rules rather than learning reusable abstractions. More broadly, generalization to harder or novel violations depends less on model scale or optimization and more on how the learning problem is structured (Chang & Bisk, 2025).
Motivated by this view, we frame code quality analysis as meta-learning (or ``meta-linting'') over code idioms, recurring patterns that capture semantic intent beyond surface syntax. In this formulation, best practices are semantic patterns over idioms, and violations occur when problematic idioms appear in contexts where safer or more appropriate alternatives exist (see Appendix C). We operationalize this by using linters to generate synthetic supervision for idioms, then train LLMs to detect and locate violations from high-level natural language specifications and illustrative examples (see Figure 1). Code idioms provide a natural intermediate abstraction, bridging simple syntactic cues and full program semantics. This allows models to leverage simpler idioms to support detection of more complex, context-dependent best practices. For example, in Python, generating passwords or tokens using random.choice is common but insecure in security-sensitive contexts. PEP 506 (D'Aprano, 2017) recommends using secrets.choice instead. Detecting such violations requires conditional reasoning: not every random.choice is problematic, but its use in security-critical code constitutes a violation. Our approach trains models to recognize both the idiom (random.choice) and the relevant context, allowing them to flag violations that go beyond simple syntactic patterns.
This highlights a broader learning challenge. Some best practices are relatively easy to detect because they rely on surface-level cues. For example, existing linters such as Ruff include rules (e.g., S104–S108) that flag likely hard-coded secrets using heuristic signals like credential-shaped strings or suspicious variable names. While limited and incomplete, these rules expose recurring cues associated with security-sensitive code. Meanwhile, hard best practices, such as those captured by PEP 506, build on these same cues but require additional semantic and contextual reasoning. Reliably detecting insecure randomness involves recognizing when such signals co-occur with developer intent, such as credential generation or authentication logic, going beyond what fixed heuristics can capture.
This defines a difficulty progression in code quality analysis driven by the increasing contextual and semantic reasoning required to recognize best practice violations. As observed by Chang & Bisk (2025), failures to generalize along such progressions stem not from data scarcity, model scale, or inference limits, but from learning paradigms that fix the hypothesis class and inhibit abstraction. In code quality analysis, flat supervision maps programs directly to violation labels, forcing models to memorize isolated best practices rather than extract reusable structure. Although easy violations expose low-level heuristics, current training setups do not allow these signals to be elevated into higher-order representations, preventing inductive transfer to harder or novel cases. In contrast, MetaLint restructures supervision around code idioms, using signals from easy violations to learn semantic abstractions that become the basis for detecting harder best practice violations from high-level specifications.
MetaLint integrates naturally with existing instruction fine-tuning (IFT) and preference-optimization (PO) methods, using them for exploiting task structure to achieve cross-task generalization (Mishra et al., 2021a; Sanh et al., 2021; Wang et al., 2022). Existing linters (e.g., Ruff (ruf, ) for Python and PMD (pmd, ) for Java) provide large-scale synthetic data by enforcing easy best practices, which we use both for supervised IFT and as verifiers during PO to improve performance on hard best practices. To systematically study this generalization, we manually construct a benchmark of hard best practices derived from popular PEPs requiring contextual reasoning. We evaluate state-of-the-art reasoning and code models on this benchmark and compare them with MetaLint trained models, examining whether they can move beyond memorizing easy heuristics.
Our key contributions are:
  1. 1.
    We introduce MetaLint, a training framework that reorganizes supervision to bootstrap learning signals for easy best practices from linters to support inductive transfer to hard and novel best practices.
  2. 2.
    We construct a benchmark of hard-to-detect, broadly relevant best practices inspired by PEPs to probe whether MetaLint models exhibit easy-to-hard generalization.
  3. 3.
    We benchmark state-of-the-art code and reasoning models on our PEP hard best practice benchmark and compare them against MetaLint-trained models. With our training Qwen3-4B outperforms larger models like o3-mini achieving the highest detection recall and comparable localization scores without test-time compute.
  4. 4.
    We show that MetaLint generalizes across programming languages (Python, Java), model families (Qwen, Llama), linters (Ruff, PMD, Tree-Sitter), reasoning settings (with and without CoT), and model scales (3B–8B).
Overall, MetaLint is not a replacement for linters, but an empirical demonstration that organizing supervision around semantic units enables LLMs to transfer learned code quality reasoning to harder or novel best practices. Finally, code quality analysis offers a realistic testbed for studying generalization: while correctness is largely objective, the difficulty progression spans from cases verifiable by automated linters to harder, context-dependent cases requiring human judgment (PEP benchmark). This work takes a first step toward such settings, where supervision becomes increasingly incomplete along the difficulty progression.
2 Related Work
Code Quality Analysis with Large Language Models. A large body of prior work has explored using LLMs for code quality analysis by interfacing them with static analysis tools or fixed rule sets. Travis Fischer (2024); lin (2023) treat LLMs as rule-guided linters via prompting or fine-tuning. While Blyth et al. (2025) proposes a static analysis-driven prompting framework to improve LLM-generated code, Du et al. (2025) conversely uses LLMs to enhance static analysis tools by reducing false-positives. Fang et al. (2025); Shin et al. (2025); Khare et al. (2023) leverage LLMs for linting and show they can outperform traditional static analysis tools. Vijayvergiya et al. (2024) train LLMs for best practice violation detection and localization, while Rasheed et al. (2024) design a multi-agent review pipeline for code review. Other works (Jiang et al., 2025b; Yao et al., 2025) use prefix-tuning and reinforcement learning with static analysis–based rewards for higher-quality, functionally correct code generation. Naik et al. (2024); Kapadnis et al. (2025); Jaoua et al. (2025) provide linter results to LLMs for more informative code reviews. Zhang et al. (2024c, b) explore AST rewrite rules and hybrid approaches combining LLMs and rules for code quality refactoring. Finally, CoUpJava (Jiang et al., 2025a) propose a Java version upgrade benchmark, mined using AST refactoring rules. Across these approaches, static analysis tools and handcrafted rules typically act as either the final decision mechanism or the primary training signal, anchoring model behavior to a fixed set of best practices. In contrast, we use static analysis as supervision for the easy end of a difficulty progression, and train models to detect best practice violations from high-level descriptions, enabling transfer to harder and previously unseen cases that extend beyond linter coverage.
Easy-to-Hard Generalization. Recent work suggests that easy-to-hard generalization arises not merely from curriculum ordering, but from how supervision is structured to promote abstraction across related tasks (Chang & Bisk, 2025). In mathematical reasoning, models trained on easier problems (e.g., levels 1–3) consistently generalize better to harder benchmarks (e.g., levels 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025), with several studies emphasizing the role of high-quality supervision, particularly for harder instances (He et al., 2024). Sun et al. (2024) show that reward models trained on simple code and math problems improve performance on more complex tasks. More broadly, (Hu et al., 2025; Gaunt et al., 2016) demonstrate that training on structurally simpler instances can yield robustness to longer or more complex reasoning problems, including code. Similarly, Zhang et al. (2024a) evaluate reward models on algorithmic tasks such as string manipulation and observe transfer from simpler to more complex formats. Most closely, Chang & Bisk (2025) argue that easy-to-hard generalization is primarily driven by the learning paradigm and finding the right hypothesis class, rather than data scarcity, model scale or inference limits. While prior work studies this phenomenon in domains such as counting, mathematics, or contest coding, code quality analysis offers a complementary setting where correctness is largely objective but supervision becomes increasingly incomplete along the difficulty progression. Additional related work on leveraging instruction following for generalization is discussed in Appendix D.
3 Method
We design the MetaLint framework to reorganize supervision for best practice violation detection, shifting the hypothesis class from memorizing fixed rules to identifying violations from high-level natural language descriptions of problematic code idioms, in line with recent work showing that such reorganization is critical for easy-to-hard generalization (Chang & Bisk, 2025). We formulate best practice violation detection as an instruction-following meta-task $M_{I}$, where for a given problematic idiom $I$, the prompt includes a natural language description $D_{I}$ and illustrative examples $E_{I}$, denoted as $M_{I}=\{D_{I},E_{I}\}$. The LLM must identify all and only those code fragments that match idiom $I$ while performing $M_{I}$. This setup discourages rote memorization by construction, as correctness is defined relative to the provided specification, and encourages adaptive reasoning over the prompt’s description. For example flagging best practice violations of any other idiom $I^{\prime}\neq I$ is penalized during $M_{I}$. By framing best practices as meta-tasks, this approach supports the LLM in remaining flexible and better aligned with evolving best practices. We describe the components of our training framework in Figure 1 and Figure 2 below.
3.1 Synthetic Data Generation
One of the main goals of our meta-task formulation is to change the learning task to detecting best practice violatons from high level specifications to encourage transfer from easy to hard cases. We train LLMs on a set of ``easy'' best practice idioms $I_{\mathcal{L}}$ that are detectable by existing linters $\mathcal{L}$, and evaluate them on a harder set $I_{\mathcal{L^{\prime}}}$ consisting of idioms that linters cannot detect (where $\mathcal{L^{\prime}}$ denotes the complement of $\mathcal{L}$, i.e., all idioms not detectable by a linter). Our hypothesis is that training on $I_{\mathcal{L}}$ exposes the model to a structured family of best practice idiom specifications, which supports inductive transfer to hard to detect best practice idioms in $I_{\mathcal{L^{\prime}}}$. Since idioms in $I_{\mathcal{L}}$ are already covered by linters, we can leverage these tools to generate large-scale synthetic training data and provide supervision. For Python, we use the popular Ruff linter, which implements over 800 rules spanning syntax modernization, security, readability, etc., while for Java, we use the PMD static analyzer, which covers 269 idioms as well as some manually written tree-sitter [^1] https://tree-sitter.github.io/tree-sitter/ queries inspired by 8 Java Enhancement Protocols (JEPs) (Table 8). We run Ruff, PMD, and the JEP tree-sitter queries on Python and Java source code files $f\in\mathcal{F}$ from the STACK (Lozhkov et al., 2024) dataset, which contains code from a diverse range of GitHub repositories. This allows us to collect files with either no violations or one or more violations for each idiom in $I_{\mathcal{L}}$. Ruff also incorporates rules from other linters such as PyFlakes, Bandit, and autoPEP8, making it well-suited for producing diverse and representative synthetic data. Additionally, to automatically build the meta-task instruction prompts $M_{I_{\mathcal{L}}}$ for each idiom, we scrape rule-specific documentation from the Ruff and PMD websites, including descriptions and examples. An example prompt, along with a code file containing lines that violate the idiom, is shown in Appendix E.1. For the JEP tree-sitter queries, since they are few in number, we manually write the meta-task prompts.
3.2 Instruction Supervised Fine-Tuning
Instruction fine-tuning serves to align the model's outputs with the natural language best practice specifications using the linter outputs as a reference for easy cases. We train the target LLM $\Phi$ on a set of linter-detectable, easy best practice idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model's output is a list of best practice violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix E.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. This formulation explicitly penalizes hallucinated violations of best practices not in the prompt, encouraging the model to flag only violations of the provided best practice and discouraging memorization. We try to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data. This results in a total of 53k synthetic training instances spanning 50 idioms (a subset of all the idioms detectable by Ruff) for Python Ruff data and 96.8k instances spanning 269 idioms for Java PMD and 127.3k instances spanning 15 idioms for tree-sitter data, respectively.
3.3 Verifiable Reward Model and Preference Optimization
We use preference optimization to further improve consistent application of the best practice specification by contrasting accurately localized violations with inaccurate ones, as determined by linter results. We adopt the RS-DPO approach (Khaki et al., 2024), which combines rejection sampling (RS) (Touvron et al., 2023) with Direct Preference Optimization (DPO) (Rafailov et al., 2023) to generate on-policy data from a supervised fine-tuned (SFT) policy model. It samples $k$ outputs per input, computes rewards for them, and constructs contrastive win–loss pairs based on the reward distribution and a threshold $\eta$ (Figure 2). We detail the verifiable linter-based reward model and contrastive pair sampling procedure below.
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter's line numbers (blue circle in ``Verifiable Reward Model'', Figure 1) as ground truth and the model’s predicted lines (yellow circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single best practice idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{w}=y_{i}$, $y_{l}=y_{j}$, and store the instance $(x,y_{w},y_{l})\in\mathcal{D}_{p}$. By constructing contrastive pairs within the same meta-task, RS-DPO encourages the model to apply a given specification consistently across variations in code structure, rather than across best practices. Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
$\Phi^{RL}=\arg\max\sum_{(x,y_{w},y_{l})\in\mathcal{D}_{p}}\log\sigma\Big($ $\beta\log\frac{\Phi^{RL}(y_{w}\mid x)}{\Phi^{SFT}(y_{w}\mid x)}$
$-\beta\log\frac{\Phi^{RL}(y_{l}\mid x)}{\Phi^{SFT}(y_{l}\mid x)}\Big)$
Here, $\sigma$ denotes the sigmoid function, and $\beta=0.1$ is the KL penalty coefficient, corresponding to low-to-moderate regularization.
3.4 Training with Reasoning Traces
Reasoning traces can act as explicit intermediate representations to support transfer across best practice idioms. To obtain CoT traces that guide the LLM to correct answers, we adopt a rejection sampling approach (``Rejection Sampling SFT'' in Figure 2) for SFT data collection. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained CoT-capable LLM (e.g., Qwen3-4B), and compute a reward $r_{y_{i}}$ for each, following the RS-DPO procedure in Figure 2. Instead of forming contrastive pairs, we discard any $y_{i}$ with $r_{y_{i}}<\gamma$, where $\gamma=1$, i.e., CoT–response pairs that are incorrect or improperly formatted. Rewards are applied only to the final response, obtained after parsing the CoT trace, and we also remove cases where the CoT fails to terminate or yield an answer. If no valid $y_{i}$ is found for an input $x$, we skip it. To promote meta-task diversity, we retain at most two valid responses per input: multiple $y_{i}$ only for violation cases and a single $y_{i}$ otherwise. This maintains the 71:29 no-violation-to-violation ratio of Ruff Python SFT data, with the latter more likely to fit within token limits. When excess valid responses exist, we keep the shortest completions, as they typically reflect more concise reasoning (final answers are of similar token length across samples). Following this policy, we collect 52.7k Python training instances from Ruff data, which we use to train the reasoning-enabled base Qwen3-4B with SFT. This yields a CoT-capable SFT model $\Phi_{CoT}^{SFT}$ for Python code quality analysis. We then apply the RS-DPO procedure in Section 3.3 and Figure 2, with the only change being that each $y_{i}$ now includes both the CoT trace and final response.
Figure 1: MetaLint: (1) Synthetic data generation with linters/tools, (2) Supervised Instruction Fine-Tuning (SFT) on this data, and (3) Verifiable Reward Model derived from the linter.
Figure 2: MetaLint: Preference Optimization using reward model: (4) Rejection Sampling Direct Preference Optimization (RS-DPO), and (5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
4 Experiments
4.1 Evaluation Metrics
We evaluate the LLM's ability to detect best practice violations through two tasks: detection, which assesses whether a best practice is violated at least once in a code file, and localization, which evaluates whether the model accurately identifies the lines with the problematic idiom. For both tasks, we report precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each best practice idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers. To handle potential class imbalance, we use macro-averaging across idioms and exclude NO VIOLATION as a class to penalize models that only predict NO VIOLATIONS FOUND (such models will score zero on all detection metrics). For localization, metrics are averaged only across instances with at least one violation in the ground truth. Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix F.1.
4.2 Generalization across Easy Best Practices
To evaluate whether MetaLint training encourages transfer to novel easy-to-detect best practices at test time, we explore transfer settings spanning Python & Java.
Ruff Python Idioms: We construct a 5.3k-instance synthetic test set spanning 50 Ruff idioms, using the data generation procedure from Section 3.1. The data has a 74:26 no-violation-to-violation split, similar to the SFT training set. Best practice idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
In domain. 5 idioms seen during training, to assess whether MetaLint improves performance on in-domain idioms.
Near transfer. 10 idioms with specifications similar but not identical to training idioms to probe memorization. Reliance on memorized patterns, may hurt performance due to interference.
Far transfer. 35 idioms distinct from training, to test whether the LLM can follow the provided specification and adapt to novel idioms at test time.
For these experiments, we use Qwen3-4B (with and without reasoning) and Llama-3.2-3B-Instruct to study the effect of test-time compute and model family.
PMD and JEP Tree-Sitter Idioms: For Java, we construct two synthetic test sets: 5.1k instances (54:46 split) spanning 269 PMD idioms, and 6.4k instances (50:50 split) spanning 15 JEP idioms (Table 5), flagged via tree-sitter queries. We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer from PMD to JEP and vice-versa. These experiments use Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct to assess the effect of model scale.
4.3 PEP Hard Best Practice Benchmark
To test whether MetaLint trained LLMs exhibit easy-to-hard generalization at test time, we construct a benchmark of hard-to-detect best practices from 15 PEPs that require abstract, context-dependent reasoning. We design heuristics per PEP (Table 15, 16 and 17) to find candidate violations in the STACK-V2 corpus, prioritizing recall to retrieve a broad set for manual selection. These best practices violations cannot be reliably detected by rule-based pattern matching, making them ideal for evaluating a model's contextual reasoning capabilities. From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (``start'' and ``end'') of the violated code, providing ground-truth for localization. We add negative examples for each PEP by picking files retrieved for a different PEP and making sure the current PEP is not violated, in order to have a balanced distribution of violation and no-violation cases. The final benchmark contains 536 examples (52% violation, 48% no-violation), enabling evaluation of easy-to-hard generalization by comparing the base model, against the SFT, and DPO-trained models. We also compare MetaLint models trained on Ruff best practices against state-of-the-art open and closed-source code and reasoning LLMs. Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B (Agarwal et al., 2025). Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 (OpenAI, 2025b). We select these models for their strong coding and reasoning performance and also evaluate the effects of code-specific pre-training, model scale (3B–120B), and test-time compute for open-source models.
5 Results
We evaluate the effectiveness of MetaLint along three dimensions: (1) generalization to unseen easy best practices (Section 5.1); (2) easy-to-hard generalization across best practices (Section 5.2); and (3) whether MetaLint enables a 4B model to achieve performance competitive with state-of-the-art code and reasoning models (Section 5.3).
5.1 Generalization across Easy Best Practices
Python Ruff Idioms: The performance of Qwen3-4B with and without reasoning and Llama3.2-3B-Instruct when trained on synthetic Ruff idioms and evaluated on the Ruff synthetic test set with varying transfer settings (section 4.2) is shown in Table 1 (full results in Table 21). While Table 1 shows the overall performance, we also analyze the performance broken down by each transfer setting in Table 18. The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop for Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics with a slight drop in detection precision. We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix F.3. Additionally, Table 18 shows that while SFT can lead to slight gains for the transfer settings, most gains emerge in the DPO stage, especially for non-reasoning models and detection recall. Overall this suggests that SFT is prone to memorizing best practices used for training, while DPO generalizes to novel best practices.
PMD and JEP Tree-Sitter Idioms: To evaluate the generality of MetaLint training across programming languages and linters, we present results from training on PMD and JEP Tree-Sitter synthetic data in Table 2 (full results in Table 31). Training on PMD shows the same overall pattern as before but with larger recall gains for both SFT and DPO, and notably stronger localization under DPO. For Llama3.1‑8B‑Instruct, SFT initially reduces detection precision, which DPO then recovers; the same precision dip‑and‑recovery appears when transferring PMD→JEP for Llama3.2‑3B‑Instruct. Despite never seeing JEP idioms during training, DPO models achieve strong detection and localization on JEP. In the untrained setting, Llama3.2‑3B‑Instruct (on PMD) and Llama3.1‑3B‑Instruct (on JEP) nearly always output the correct format but predict NO VIOLATIONS FOUND, yielding zero or near‑zero scores, since our metrics exclude that class for detection and only score positive cases for localization. Training on JEP yields high in‑domain performance for all metrics with minimal additional benefit from DPO, likely due to JEP’s smaller idiom set (15 vs 269 for PMD) and more precise instructions (Table 5). In the harder JEP→PMD transfer, DPO outperforms SFT, though overall transfer remains weaker than PMD→JEP, reflecting PMD's broader diversity and more challenging specifications (Appendix E.5).
Overall, MetaLint outperforms the base model on novel best practice violations, though gains depend on the diversity of training idioms and the gap in instruction quality between training and test data.
Table 1: Cross-Idiom Generalization on Python Ruff Idioms: Effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning). Best score across the compared training setups per model are bolded.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5380 0.2637 0.3539 0.1396 0.1479 0.1436
Qwen3-4B + SFT 0.7686 0.3178 0.4497 0.2976 0.2960 0.2968
Qwen3-4B + SFT + RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
Qwen3-4B w CoT 0.8812 0.6854 0.7710 0.5049 0.4878 0.4962
Qwen3-4B w CoT + RS-SFT 0.9350 0.8183 0.8727 0.6639 0.6500 0.6569
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.7710 0.7571 0.7640
Table 2: Cross-Idiom Generalization on JEP & PMD Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 31). The transfer column indicates training and test data on the left and right side of the arrow. Best score across the compared training setups per model are bolded.
Model Transfer Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Llama3.2-3B-Instruct PMD $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
Llama3.2-3B-Instruct + SFT 0.2251 0.4421 0.2983 0.2822 0.2778 0.2800
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4395 0.8908 0.5886 0.5930 0.5969 0.5949
Llama3.2-3B-Instruct PMD $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
Llama3.2-3B-Instruct + SFT 0.2286 0.4072 0.2928 0.1626 0.1336 0.1467
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
5.2 Evaluating Easy-to-Hard Generalization
To evaluate whether MetaLint training on easy to detect Ruff best practices improves performance on hard, PEP best practices, we report results on our PEP hard best practice benchmark (Table 3, full results in Table 22). At the SFT stage, performance declines for Qwen3-4B (with and without CoT) but improves slightly for Llama3.2-3B-Instruct, suggesting that SFT is prone to memorization of the training distribution. In contrast, DPO yields clear improvements in detection and localization (except detection precision for Llama3.2-3B-Instruct), with statistically significant gains (Appendix G.2). An additional experiment training Qwen3-4B (CoT) directly with RS-DPO, bypassing SFT, resulted in near-zero performance because many generated DPO pairs violated the required output format, which the model inherited. Thus, SFT, despite its drawbacks, is essential for teaching format compliance and setting the stage for DPO to encourage easy-to-hard generalization. Interestingly, the non-CoT model achieves substantially higher detection recall and slightly higher F-score than the CoT variant, despite lower precision. Our analysis attributes the CoT model's reduced recall to its more conservative interpretation of idiom specifications and to errors such as misinterpretation, overthinking, and skipped lines, as detailed in Appendix G.3. Finally we analyze the impact of the illustrative examples in the prompt in Table 19.
Table 3: Easy-to-Hard Generalization: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B + SFT 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
Qwen3-4B + SFT + RS-DPO 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B w CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-4B w CoT + RS-SFT 0.7615 0.3689 0.4970 0.2785 0.1437 0.1896
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Table 4: Benchmarking on Hard Best Practices: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training. The best scores are bolded and second best and underlined.
Detection Localization
Model $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
Qwen3-14B 0.9021 0.4612 0.6103 0.2890 0.2521 0.2693
Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.3540
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.1630 0.1477 0.1550
Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.7130 0.4015 0.3403 0.3684
GPT-oss-20b 0.8377 0.3531 0.4968 0.2510 0.1695 0.2024
GPT-oss-120b 0.9157 0.6456 0.7573 0.3991 0.3331 0.3631
Qwen3-4B MetaLint (SFT+RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B MetaLint w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.3320 0.3807
GPT-4.1 0.9070 0.6460 0.7546 0.4632 0.4673 0.4653
GPT-5 (high) 0.9130 0.5673 0.6998 0.4397 0.4257 0.4326
5.3 Benchmarking on Hard Best Practices
Table 4 compares the best-performing Qwen3-4B MetaLint DPO models against state-of-the-art code and reasoning models (full results in Table 20).
Detection: For detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini). However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix G.2) and outperform GPT-oss-20B. This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy best practices, and that the non-CoT model does not use test-time compute. We also identify localization errors and their frequency in Table 32.
Overall, our models perform comparably to state-of-the-art systems, with the non-CoT Qwen3-4B variant achieving the highest recall, and both MetaLint-trained Qwen and LLaMA models exhibiting evidence of easy-to-hard generalization. This aligns with Chang & Bisk (2025), which highlights the role of reorganizing supervision and learning paradigm, rather than scaling or optimization alone, in enabling generalization along difficulty progressions.
6 Conclusion and Future Work
Our results show that MetaLint encourages inductive transfer over best practice specifications, rather than memorization of fixed rules. We observe generalization to unseen best practices across Python and Java, multiple linters, model families, reasoning settings, and model scales. Models trained with MetaLint also exhibit easy-to-hard generalization, transferring from linter-detectable best practices to harder PEP violations. With only 4B parameters, MetaLint-trained Qwen models achieve detection performance comparable to strong code and reasoning systems, achieving the highest recall and competitive precision. Localization remains more challenging but is competitive with o3-mini and surpasses GPT-oss-20B without test-time compute. Overall, these results support the view that reorganizing supervision, rather than scaling alone, enables inductive generalization beyond the training distribution (Chang & Bisk, 2025). MetaLint is not a replacement for static analysis, but a proof of concept that inductive learning can improve LLM performance along a natural difficulty progression in code quality analysis. Future work will target automated refactoring and better RL training like Group Relative Policy Optimization (GRPO) (Shao et al., 2024).
Acknowledgements
This work was partially supported by a generous collaborative research grant from Oracle Labs. We are grateful to Jack Sullivan, Qinlan Shen, Adam Pocock, and Nidhi Chandra at Oracle Labs for their guidance, feedback, and insightful discussions that played a key role in shaping the core ideas behind MetaLint and advancing the project.
Impact Statement
This work aims to improve the reliability and generalization of automated code quality analysis by enabling models to adapt to evolving and previously unseen best practices. By reducing reliance on brittle, static rule sets, MetaLint has the potential to lower maintenance costs for software tooling and improve developer productivity and security outcomes, particularly in settings where best practices change rapidly. However, as with any automated code analysis system, incorrect or overconfident recommendations could mislead developers or create false assurance if deployed without human oversight. We mitigate this risk by grounding supervision in existing static analysis signals and framing outputs as assistive diagnostics rather than authoritative judgments. We do not anticipate significant negative societal impacts beyond those common to developer tooling, and we expect the primary effects of this work to be beneficial.
Appendix A Appendix
We organize the Appendix in five parts: 1) Limitations of our work (Appendix B), 2) Clarifications and definitions of terms used in the paper (Appendix C), 3) Additional related work about instruction following and generalization (Appendix D), 4) Method details like prompts and hyperparameters (Appendix E), 5) Additiomal experimental details (Appendix F) and 6) Additional results and ablations (Appendix G).
Appendix B Limitations
Despite promising results, including evidence of easy-to-hard generalization, MetaLint has several limitations that suggest directions for future work:
  • Chain-of-Thought (CoT) reasoning: We did not explore whether non-CoT models can be trained to generate CoT-style reasoning using supervision from a teacher model.
  • Self-improvement and data generation: We attempted self-improvement strategies for RS-SFT data generation (e.g., STaR (Zelikman et al., 2022)) when the base model failed. However, generating CoTs that do not directly reference provided hints proved challenging and risks contaminating the training data. We therefore adopted a simpler rejection sampling (RS-SFT) strategy.
  • Preference optimization methods: We only experimented with DPO for preference optimization. Alternative approaches, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (Shao et al., 2024), combined with our verifiable linter-based reward model, could be explored.
  • Difficulty progression and curriculum: Our current setup does not implement a fine-grained difficulty progression or a curriculum over meta-task abstractions. Incorporating such structure could encourage more systematic induction over code idioms and code quality concepts.
  • Multilingual training: Experiments focus on one programming language at a time (Python or Java). Future work could explore joint training across multiple languages, cross-language transfer, and structured learning that encourages abstraction across similar language constructs (e.g., Python’s PEP 557: Data Classes vs. Java’s JEP 395: Records).
  • Task scope: We focus exclusively on best practice violation detection and do not yet explore inductive transfer over code refactoring or other code quality tasks.
Appendix C Clarification and Term Definitions
Because this paper studies a software engineering–heavy problem setting, we clarify the key technical terms used throughout.
Code Idiom / Programming Idiom.
A code idiom is a recurring, language-specific pattern of code that conveys a particular semantic intent or programming practice beyond what is expressed by its literal syntax.
Code Quality Analysis.
Code quality analysis refers to the systematic evaluation of source code with respect to software engineering attributes such as readability, complexity, maintainability, and adherence to best practices, using automated tools or model-based methods. Although best practice analysis is only one component of code quality analysis, we use these terms interchangeably in this paper for brevity.
Best Practice Violation Detection.
Best practice violation detection is a subproblem of code quality analysis that involves identifying and localizing code fragments that violate established or recommended programming practices. In this work, we operationalize best practice violation detection as the task of identifying problematic code idioms, i.e., recurring patterns that correspond to non-recommended practices. While this formulation is more restrictive than the full space of best practices, it enables a systematic task structure that supports controlled supervision and generalization. Additionally we often use the shorthand best practice idiom or idiom to refer to the problematic idiom that needs to be detected to flag a best practice violation.
Appendix D More Related Work
Instruction Following and Generalization. Instruction tuning has emerged as a powerful form of meta-learning, enabling cross-task generalization by training models to interpret and follow natural language instructions rather than memorizing fixed tasks. Prior work shows that exposure to diverse task instructions allows models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, substituting for explicit supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity drives generalization: varied instructions outperform repeated exposure to identical formats (Charton et al., 2024). This effect holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness but relying entirely on LLM-generated data without verifiers. These results suggest that instruction tuning functions as task-level meta-learning, enabling models to adapt to new tasks through natural language. Building on this insight, we model individual code quality idioms as separate meta-tasks and generate large-scale synthetic data for each task to support cross-idiom generalization. This approach allows the trained model to adapt to new idioms and evolving best practices.
Appendix E Method Additional Details
E.1 MetaLint Instruction Following Prompt
We used the following instruction following style prompt to train the model with synthetic Ruff best practice idiom data for the meta-linting task:
MetaLint Instruction Following Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} Violations per idiom:
An example input with the code file and best practice idiom spec populated as well as the expected JSON style output is shown below:
Example Ruff Meta-Task Input Look at the following list of code idiom specifications with definitions and examples: # Idiom ANN202 (missing-return-type-private-function) Definition: Checks that private functions and methods have return type annotations. Rationale: Type annotations are a good way to document the return types of functions. They also help catch bugs, when used alongside a type checker, by ensuring that the types of any returned values, and the types expected by callers, match expectation. Example:
def _add(a, b):
    return a + b
Use instead:
def _add(a: int, b: int) -> int:
    return a + b
Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say 'NO VIOLATIONS FOUND' if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file:
  1 # -*- coding: utf-8 -*-
  2 # pragma pylint: disable=unused-argument, no-self-use
 86     def _reload(self, event, opts):
 87         """Configuration options have changed,
    save new values"""
 88         self.options = opts.get("fn_cisco_amp4ep", {})
 89         validate_opts(self)
 90
 91     @function("fn_amp_move_computer")
 92     def _fn_amp_move_computer_function(self, event, *args,
    **kwargs):
 93         """Function: Move computer to a group with given
    connector guid and group guid."""
 94         try:
Violations per idiom:
Example Ruff Meta-Task Output **Idiom ANN202 Violations:**
{"line": " 86     def _reload(self, event, opts):", "fix": null}
{"line": " 92     def _fn_amp_move_computer_function(self,
event, *args, **kwargs):", "fix": null}
E.2 DPO Contrastive Pair and RS-SFT Sampling Details
To generate RS-DPO contrastive samples (or RS-SFT outputs) from the baseline SFT (or untrained) models, we used the following hyperparameters: nucleus sampling with a maximum of 2048 new tokens, $k=5$ sampled outputs per input, temperatures picked cyclically from {0, 0.3, 0.5, 0.7, 1}, a top-$p$ (cumulative probability threshold) of 0.95, and a seed of $42+i$, where $i\in\{1,\dots,k\}$, to encourage both reproducibility and output diversity.
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section E.4.
E.3 Training Hyperparameters and Computational Environment
Python SFT/RS-SFT hyperparameters:
We fine-tune the Qwen3-4B model using flash_attention_2 and bfloat16 precision. The model is trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and a warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled to reduce memory usage, with non-reentrant mode. Evaluation is performed every 2000 steps, and checkpoints are saved at the same interval. Special tokens are manually handled in the chat template without automatic insertion. The training uses 12 preprocessing workers and is seeded with 42 for reproducibility.
Python RS-DPO parameters:
We fine-tune the model using RS-DPO with bfloat16 precision and a reward shaping parameter $\beta=0.1$. Training is performed for 1 epoch with a learning rate of 5e-7, cosine learning rate scheduling, and a warmup ratio of 0.1. We use a maximum input length of 3500 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing is enabled with non-reentrant mode to optimize memory usage. The optimizer is AdamW, and evaluation is conducted every 200 steps with checkpoints saved at the same interval. The training is seeded with 42 for reproducibility.
Java SFT hyperparameters:
For Java experiments, we fine-tune Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct with bfloat16 precision. Both models are trained for 2 epochs with a learning rate of 2e-5, cosine learning rate schedule, and warmup ratio of 0.1. We use a maximum sequence length of 3000 tokens, per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing (non-reentrant) is enabled. Evaluation and checkpoint saving occur every 5000 steps. Special tokens are manually handled in the chat template. Training is seeded with 42.
Java RS-DPO parameters:
RS-DPO training is performed on Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct using bfloat16 precision. Training runs for 1 epoch with a learning rate of 5e-7, cosine learning rate scheduling, and warmup ratio of 0.1. We use a maximum input length of 3500 tokens, a per-device batch size of 2, and gradient accumulation steps of 4. Gradient checkpointing (non-reentrant) is enabled. Evaluation and checkpoints are recorded every 200 steps. Reward shaping parameters vary across settings, with $\beta\in\{0.1,0.5,1\}$. Seeds are fixed at 42 for reproducibility.
Computational Environment:
All SFT, RS-SFT, and RS-DPO experiments (Python and Java) were conducted on a Linux server equipped with NVIDIA A100 80GB GPUs (Ampere architecture), CUDA 12.9, and driver version 575.51.03. Each job had access to 100 GB of CPU memory and 2 CPU cores. Training used mixed-precision (bfloat16) with gradient checkpointing to optimize memory usage. Inference used a similar setup with GPU allocation varying by model size.
E.4 Baseline Inference Details
We use the following hyperparameters for performing inference with the baseline LLMs:
Open Source LLMs: We perform nucleus sampling with 8192 max-new tokens, temperature of 0.7, top-p (cumulative probability threshold) of 0.95 and seed of 42 (to promote reproducibility).
Closed Source LLMs: We use the chat completion OpenAI API with max tokens of 1024 for GPT-4.1 and GPT-4o and max completion tokens of 3000 for o3-mini and o4-mini. We use default parameters for everything else (temperature of 1 and top-p of 1, no presence penalty). For GPT-5 we use 8192 max completion tokens and high reasoning effort.
Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches. For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Best Practice Benchmark.
Baseline Inference Prompt Look at the following list of code idiom specifications with definitions and examples: {LIST_OF_IDIOM_SPECS} Given these idioms, your task is to look at a code file and detect violations of the above idioms, and flag them like a linter. You should also suggest a fix if possible. Report the results per idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for a given idiom. Do not detect any idioms not specified above. Code file: {CODE_FILE} # OUTPUT FORMAT I want you to generate your output under a section called ``### Final Idiom Violations Found''. Structure you response for a given idiom XYZ as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:**
[…]
Inquiry.objects.filter(owner=kwargs.pop('user')))", "fix": null}
[…]
4 from inquiry.models import *",
[…]
and as follows for cases with violations: ### Final Idiom Violations Found **Idiom XYZ Violations:** NO VIOLATIONS FOUND Violations per idiom:
Table 5: JEP Best Practice Idiom Specifications (1/3): This table presents 15 best practice idioms across 8 JEPs, including both ``before'' (old best practice) and ``after'' (updated best practice) patterns. The JEP# column lists the JEP number, the JEP title specifies the idiom topic, and the parenthesized value indicates whether it is a before or after pattern. The Definition, Example, and Tree-Sitter Queries columns provide the idiom definition, minimal Java examples shown to the LLM as instructions, and the queries used to flag problematic idioms for synthetic data creation.
JEP# JEP Title Definition Example(s) Tree Sitter Queries
394
PatternMatching
InstanceOf
Usage of the old pattern
of testing with instanceof
followed by a manual cast
to extract and operate on
the object. This pattern is
verbose and repetitive. Flag
the instanceof expression
check within a conditional
statement and the
accompanying cast
expression in the body of
the conditional statement.
public class ShapeExample {
static double getPerimeter(Object obj) {
if (obj instanceof Rectangle) {
Rectangle r = (Rectangle) obj;
return 2 * r.length() + 2 * r.width();
} else if (obj instanceof Circle) {
Circle c = (Circle) obj;
return 2 * c.radius() * Math.PI;
} else {
throw new IllegalArgumentException(
"Unrecognized shape");
(if_statement
condition: (parenthesized_expression
(instanceof_expression
left: (identifier) @H1
right: (type_identifier) @H2
) @jep_394_before_instanceof_expression.part1
consequence: (block
((local_variable_declaration
type: (type_identifier) @H3
declarator: (variable_declarator
value: (cast_expression
type: (type_identifier) @H4
value: (identifier) @H5
)(#eq? @H1 @H5) (#eq? @H2 @H3) (
#eq? @H3 @H4)
) @jep_394_before_instanceof_expression.part2
394
PatternMatching
InstanceOf
Replaces verbose instanceof
tests plus manual casting into
a concise form that tests and
declares a typed variable in
one step, for example,
"if (obj instanceof String s)"
which improves readability,
reduces boilerplate, and
introduces flow-scoped pattern
variables. Flag only the line
containing the combined
instancesof test and casting
within the conditional
statement.
public class ShapeExample {
static double getPerimeter(Object obj) {
if (obj instanceof Rectangle r) {
return 2 * r.length() + 2 * r.width();
} else if (obj instanceof Circle c) {
return 2 * c.radius() * Math.PI;
} else {
throw new IllegalArgumentException(
"Unrecognized shape");
(instanceof_expression
left: (_)
right: (type_identifier)
name: (identifier)
) @jep_394_after_instanceof_expression
378
TextBlocks
Multiline strings represented
using concatenated string
literals, requiring explicit
newline escape sequences
[…]
with the + operator. This
approach is verbose and
error-prone. Flag cases
where a variable
declaration or method
invocation uses
concatenated string literals
instead of multiline strings.
[…]
(local_variable_declaration
declarator: (variable_declarator
name: (identifier)
value: [
(binary_expression
] @jep_378_before_concatenated_string_literals
378
TextBlocks
Use of multiline string literal
enclosed by triple
double-quote marks ("""),
allowing for cleaner and more
readable representation of
multiline strings without
explicit escape sequences.
Flag cases that use triple
double-quote marks for
multiline strings in variable
declarations or
method invocations.
String html = """
<html>
<body>
<p>Hello, world!</p>
</body>
</html>
(string_literal) @jep_378_after_text_block
(#match?
@jep_378_after_text_block "^\"\"\"")
361
Switch
Expressions
Misuse of switch statement
with fall-through behavior
'for pattern matching. This
pattern is verbose and error
prone. You should flag case
statements with empty
bodies that are misusing
fall-through behavior.
int numLetters;
switch (day) {
case MONDAY:
case FRIDAY:
case SUNDAY:
numLetters = 6;
break;
case TUESDAY:
numLetters = 7;
break;
throw new IllegalStateException(
"Unexpected value: " + day);
jep_361_before_custom_detectors
Table 6: JEP Best Practice Idiom Specifications (2/3)
JEP# JEP Title Definition Example(s) Tree Sitter Queries
361
Switch
Expressions
Use of switch expressions, allowing
a return value. Employs the ->
syntax for case labels, eliminating
fall-through behavior. Flag
statements that use the arrow
operator "->" or "yield" syntax.
Example 1:
int numLetters = switch (day) {
case MONDAY, FRIDAY, SUNDAY ->6;
case TUESDAY ->7;
case THURSDAY, SATURDAY ->8;
case WEDNESDAY ->9;
Example 7:
String category = switch (age) {
case 0, 1, 2, 3, 4, 5 ->"Toddler";
case 6, 7, 8, 9, 10, 11, 12 ->"Child";
case 13, 14, 15, 16, 17, 18, 19 ->"Teenager";
default ->"Adult";
Example 8:
String response = switch (input) {
case "yes" ->"Affirmative";
case "no" ->"Negative";
default ->"Unrecognized input";
(yield_statement
) @jep_361_after_yield
(switch_rule
(switch_label)
"->" @jep_361_after_arrow
(switch_rule
(switch_label)
"->" ;; ensures it's not arrow
(block (yield_statement
) @jep_361_after_yield)
314
UnicodeLang
TagExtensions
Use java.util.Locale with additional
BCP 47 Unicode extensions
(cu, fw, rg, tz) in Java 10 to customize
locale behavior like currency
(java.util.Currency), first-day-of-week
(java.time.temporal.WeekFields),
region override
(java.text.NumberFormat.getInstance),
and time zone
(java.time.format.DateTimeFormatter).
Flag imports and function calls
related to these.
Example 1: Currency Type (cu)
import java.util.Locale;
import java.util.Currency;
public class Foo {
void bar() {
Locale locale = Locale.forLanguageTag(
"en-US-u-cu-EUR");
Currency c = Currency.getInstance(locale);
System.out.println(c);
Example 4: Time Zone (tz)
import java.util.Locale;
import java.time.format.DateTimeFormatter;
import java.time.ZonedDateTime;
public class Foo {
void bar() {
Locale locale = Locale.forLanguageTag(
"en-US-u-tz-Asia-Tokyo");
DateTimeFormatter fmt =
DateTimeFormatter.ofPattern(
"yyyy-$M$-$d$ $H$:$m$ z").withLocale(locale);
System.out.println(
fmt.format(ZonedDateTime.now()));
(import_declaration
((scoped_identifier
scope: (scoped_identifier
) @H2
name: (identifier) @H1
) (#eq? @H2 "java.util") (
#eq? @H1 "Currency"))
) @jep_314_after_currency_import
((method_invocation
object: (identifier) @H7
name: (identifier) @H8
) (#eq? @H7 "Currency") (
#eq? @H8 "getInstance")
) @jep_314_after_currency…
((method_invocation
object: (identifier) @H13
name: (identifier) @H14
#eq? @H13 "NumberFormat") (
#eq? @H14 "getInstance")
) @jep_314_after_number_format…
395
RecordClass
Use of simple data aggregates with
traditional classes which could be
replaced with a record class. This
approach requires explicit
declarations of fields, constructors,
and accessor methods, leading to
verbose and repetitive code. Flag
non record classes containing equals(),
hashCode(), and toString() methods.
public class Point {
private final int x;
private final int y;
public int x() {
return x;
@Override
public String toString() {
return "Point{x=" + x + ", y=" + y + "}";
@Override
public boolean equals(Object obj) {
@Override
public int hashCode() {
return Objects.hash(x, y);
(class_declaration
body: (class_body
(constructor_declaration
) @H1
(method_declaration
name: (identifier) @H2
) (#match? @H2 "^(
hashCode|equals|toString)\$")
) @jep_395_before_record_like_class
395
RecordClass
Use of record class. Record classes
introduce a concise syntax for
defining immutable data aggregates,
automatically generating canonical
constructors, accessors, equals(),
hashCode(), and toString() methods,
thereby reducing boilerplate code
and enhancing readability.
Example 1 (Record Declaration):
record Point(int x, int y) {}
Example 2 (Record Declaration):
record Rectangle(double length, double width) {}
(record_declaration
) @jep_395_after_record_…
Table 7: JEP Best Practice Idiom Specifications (3/3)
JEP# JEP Title Definition Example(s) Tree Sitter Queries
409
Sealed
Class
Use of abstract classes with
private constructors to simulate
sealed classes using
package-private visibility to
restrict subclassing. This
approach lacks explicit language
support and is error-prone.
Switch to sealed classes. Flag
abstract classes with private
constructors.
public abstract class Shape {
private Shape() {}
public class Circle extends Shape {
/* Implementation */
public class Square extends Shape {
/* Implementation */
((class_declaration
(modifiers) @H1
name: (identifier) @H4
body: (class_body
(constructor_declaration
(modifiers) @H2
name: (identifier) @H3
… )(#match? @H1 "abstract")
(#eq? @H3 @H4)
(#match? @H2 "private")
) @jep_409_before_abstract_class…
409
Sealed
Class
Use of sealed classes to
explicitly define which classes
or interfaces can extend or
implement them using the
sealed modifier and the permits
clause. This feature enhances
type safety and exhaustiveness
checking. Flag class declarations
with the sealed or non-sealed
modifiers and lines with the
permit clause.
public sealed class Shape
permits Circle, Square {
/* Implementation */
public final class Circle extends Shape {
/* Implementation */
public final class Square extends Shape {
/* Implementation */
(permits
) @jep_409_after_permits_clause
((class_declaration
(modifiers) @H1
)(#match? @H1 "sealed")
) @jep_409_after_sealed_modifier
406
Pattern
Matching
Switch
Use of a sequence of if-else
if statements to test an object's
type via instanceof, with a
manual cast, to handle each case
separately. This approach is
verbose, error-prone, and lacks
exhaustiveness checking or
compiler assistance for missing
cases. Flag if or else-if statements
that contain instanceof statements
with a manual cast in the
statement body.
static String formatter(Object o) {
if (o instanceof Integer) {
Integer i = (Integer) o;
return String.format("int %d", i);
} else if (o instanceof Long) {
Long l = (Long) o;
return String.format("long %d", l);
} else if (o instanceof String) {
String s = (String) o;
return String.format("String %s", s);
} else {
return o.toString();
(if_statement
condition: (parenthesized_expression
(instanceof_expression
left: (identifier) @H1
right: (type_identifier) @H2
) @jep_406_before_if_else_if_…
consequence: (block
((local_variable_declaration
type: (type_identifier) @H3
declarator: (variable_declarator
value: (cast_expression
type: (type_identifier) @H4
value: (identifier) @H5
(#eq? @H1 @H5)
(#eq? @H2 @H3)
(#eq? @H3 @H4)
) @jep_406_before_if_else_if…
406
Pattern
Matching
Switch
Use of a switch expression or
statement with case labels
containing type patterns (and
optionally a guard), binding
the matched variable within
the branch. This style is more
concise, expressive, and opens
opportunities for compiler-checked
exhaustiveness and performance
optimizations. Flag switch labels
(case statements) with patterns, null
literals or paranthesized expressions
but skip default switch labels/cases.
Example 1:
static String formatter(Object o) {
return switch (o) {
case Integer i ->String.format("int %d", i);
case Long l ->String.format("long %d", l);
case String s ->String.format("String %s", s);
default ->o.toString();
Example 2:
static String checkShape(Object o) {
return switch (o) {
(switch_label
(null_literal)
) @jep_406_after_null_case
(switch_label
) @jep_406_after_switch_pattern
(switch_label
(parenthesized_expression)
) @jep_406_after_paranthesized_pattern
(switch_label
(binary_expression)
) @jep_406_after_binary_expression
323
LocalVar
Syntax
Lambda
Params
Use of implicitly typed lambda
expressions with omitted type
declarations. These lambda
expressions rely solely on parameter
names. This approach prioritizes
brevity but lacks explicit type
information. Flag full lambda
expressions without type
declarations.
Example 1:
xs.stream().filter((a, b) ->a <b).forEach(
System.out::println);
Example 4:
xs.stream().filter((a) ->a >10).forEach(
System.out::println);
jep_323_before_custom_detector
323
LocalVar
Syntax
Lambda
Params
Use of explicit type declarations for
lambda parameters, enhancing code
clarity and enabling better static
analysis tools. Flag full lambda
expressions with explicit type
declarations using formal
parameters (var).
Example 1:
xs.stream().filter(
(var a, var b) ->a.compareTo(b) <0).forEach(
System.out::println);
Example 4:
xs.stream().filter((var a) ->a >10).forEach(
System.out::println);
(lambda_expression
parameters: (formal_parameters
(formal_parameter
type: (type_identifier) @H1
(#eq? @H1 "var"))
)) @jep_323_after_local_var_lambda
E.5 PMD Best Practice Idiom Specifications
We scrape PMD idioms specification from the Java section of the PMD rules documentationhttps://docs.pmd-code.org/latest/pmd_rules_java.html. The PMD instructions are more complex and more ambiguous than our handcrafted JEP specifications because the examples are more verbose and don't pinpoint the specific lines that should be flagged as best practice idiom violations, as can be seen in the example below.
PMD Rule Specification: UnitTestShouldIncludeAssert
Since: PMD 2.0
Priority: Medium (3)
Unit tests should include at least one assertion. This makes
the tests more robust, and using assert with messages provide
the developer a clearer idea of what the test does. This rule
checks for JUnit (3, 4 and 5) and TestNG Tests. Note: This rule
was named JUnitTestsShouldIncludeAssert before PMD 7.7.0. This
rule is defined by the following Java class:
net.sourceforge.pmd.lang.java.rule.bestpractices.
UnitTestShouldIncludeAssertRule
Example(s):
public class Foo {
   @Test
   public void testSomething() {
      Bar b = findBar();
      // This is better than having a NullPointerException
      // assertNotNull("bar not found", b);
      b.work();
This rule has the following properties:
Name
Default Value
Description
extraAssertMethodNames
Extra valid assertion methods names
Use this rule with the default properties by just referencing
it:
<rule ref="category/java/bestpractices.xml/
UnitTestShouldIncludeAssert" />
Use this rule and customize it:
<rule ref="category/java/bestpractices.xml/
UnitTestShouldIncludeAssert">
    <properties>
        <property name="extraAssertMethodNames" value="" />
    </properties>
</rule>
Table 8: List of JEPs addressed by our tree-sitter synthetic data. The JEP# and Title column indicate the number and title of the JEP while JDK# and Release Date indicate the JDK needed for compilation to be able to use the JEP features. The Before and After columns indicate whether we include rules/patterns to flag the old problematic idiom or new recommended idiom introduced by the JEP.
JEP # Before After Title JDK# Release Date
409 Yes Yes Sealed Classes 17
406 Yes Yes Pattern Matching for switch 17 14 Sept 2021
395 Yes Yes Records 16
394 Yes Yes Pattern Matching for instanceof 16 16 Mar 2021
378 Yes Yes Text Blocks 15 15 Sept 2020
361 Yes Yes Switch Expressions 14 17 Mar 2020
323 Yes Yes
Local-Variable Syntax for
Lambda Parameters
11 25 Sept 2018
314 No Yes
Additional Unicode
Language-Tag Extensions
10 20 Mar 2018
Java MetaLint Instruction Following Prompt Task Instructions (1/2): Look at the following code idiom specification with definitions and examples: {IDIOM_SPEC} Task Instructions (2/2): Given this idiom, your task is to look at a code file and detect violations of the above idiom, and flag them like a linter. You should also suggest a fix if possible. Report the results for only the idiom specification mentioned above and just say NO VIOLATIONS FOUND if no violations are found for the given idiom. Do not detect violations of any idiom not specified above. Code file: {CODE_FILE} Violations per idiom:
Appendix F Additional Experimental Details
F.1 Evaluation Metrics
Let $I$ denote an best practice idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations. For each dataset instance with input prompt $x$ and ground truth set of line numbers $y$, $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$.
We define the indicator variable:
$\mathds{1}[x]=\begin{cases}1&\text{if }x\text{ is true}\\ 0&\text{otherwise}\end{cases}$
Detection Metrics:
$P_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|=0]\cdot\mathds{1}[|\hat{y}|>0]\right)}$
$R_{I}=\frac{\sum_{(x,y)\in\mathcal{D}}\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]}{\sum_{(x,y)\in\mathcal{D}}\left(\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|>0]+\mathds{1}[|y|>0]\cdot\mathds{1}[|\hat{y}|=0]\right)}$
Macro-averaged detection metrics:
$P_{\text{Det}}=\frac{1}{|I|}\sum_{I}P_{I},\quad R_{\text{Det}}=\frac{1}{|I|}\sum_{I}R_{I},\quad F_{\text{Det}}=\frac{2P_{\text{Det}}R_{\text{Det}}}{P_{\text{Det}}+R_{\text{Det}}}$
Localization Metrics:
$P_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|\hat{y}|},\quad R_{\text{Loc}}=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\frac{|y\cap\hat{y}|}{|y|},\quad F_{\text{Loc}}=\frac{2P_{\text{Loc}}R_{\text{Loc}}}{P_{\text{Loc}}+R_{\text{Loc}}}$
F.2 Idioms Chosen for Ruff Best Practice Transfer Dataset
Table 9 lists the Ruff best practice idioms used in the SFT training and synthetic transfer evaluation test sets. Idioms are grouped by their source linter and cover a range of syntax, semantics, naming, and upgrade-related rules.
Table 9: Ruff best practice idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
Training Set Idioms Test Set Idioms
PyFlakes: PyFlakes:
F405, F501, F502, F601, F621 F403, F406, F503, F602, F622
pycodestyle: pycodestyle:
E402, E701, E721, E741, E743 E401, E702, E722, E731, E742
Naming: Miscellaneous:
N801, N802, N803, N804, N805, ERA001, C901, I001, I002, BLE001
N806, N807, N811, N812, N813 (shared with training)
pyupgrade: flake8 annotations:
UP001, UP002, UP003, UP004, UP005, UP006, ANN001, ANN002, ANN003, ANN201, ANN202,
UP007, UP008, UP009, UP010, UP011, ANN204, ANN205, ANN206
UP040, UP044, UP045, UP046, UP047
Miscellaneous: flake8 async:
ERA001, C901, I001, I002, BLE001 ASYNC100, ASYNC105, ASYNC109, ASYNC110,
ASYNC115, ASYNC116, ASYNC210, ASYNC220,
ASYNC221, ASYNC222, ASYNC230, ASYNC251
Bugbear: flake8 bandit:
B002, B003, B004, B005, B006, S102, S103, S104, S105, S106,
B007, B008, B009, B010, B012 S107, S108, S110, S112, S113,
S201, S202, S301, S302, S303
Figure 3: ID: In-Domain, NeT: Near Transfer, FaT: Far Transfer.
F.3 DPO No Violation Fraction Ablations
We analyze the impact of varying the amount of samples with zero violations used for RS-DPO training. These experiments were motivated by initial findings comparing models trained only on data with at least one violation to those trained on the full dataset. By design, RS-DPO generates significantly more training data for cases with at least one violation, due to greater variance in reward signals. This is further amplified by the fact that the initial SFT policy/checkpoint is already quite accurate in handling cases with NO VIOLATIONS FOUND leading to low variance in reward across responses.
Our early experiments showed that excluding all NO VIOLATIONS FOUND cases led to notable gains in recall and line-level localization. However, this came at the cost of a significant drop in precision compared to the SFT policy/base model. Further analysis revealed a sharp decline in the accuracy of predicting NO VIOLATIONS FOUND, from nearly 99% down to 70-80%, with performance worsening monotonically over training steps. Conversely, training on the full dataset (i.e., including 100% of the NO VIOLATIONS FOUND cases) improved precision but offered only modest gains in recall and localization, which also degraded with continued training. These findings suggest that while some NO VIOLATIONS FOUND data is necessary to maintain high precision, too much of it may hinder recall and localization.
To investigate this trade-off, we experimented with keeping only a fraction of the NO VIOLATIONS FOUND data during training. Specifically, we randomly sampled kk% of such data, varying $k$ across {0%, 2%, 5%, 10%, 20%, 40%, 100%}. These percentages were selected based on observed trends: 20%, 40%, and 100% yielded similar results, which discouraged further tests at 60% or 80%, while 2% and 5% were chosen due to a noticeable performance jump between 0% and 10%. We found that 5% offered a favorable middle ground, largely retaining or slightly reducing precision, while preserving most of the recall (resulting in the highest detection F-score), and only modestly impacting line-level localization. Based on these insights, we conducted a limited ablation on the CoT model, evaluating 2% and 5% inclusion to determine the optimal setting for both detection and localization (as shown in Table 11).
Table 10: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT. Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND. Conversely, including 100% improves precision but leads to reduced recall and localization performance. All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Best Practice Transfer test set.
Fraction of NV data Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
0% 0.6268 0.9577 0.7577 0.6777 0.6932 0.6854
2% 0.671 0.9128 0.7734 0.6681 0.6812 0.6746
5% 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
10% 0.7584 0.8114 0.784 0.6263 0.6474 0.6367
20% 0.8382 0.7227 0.7762 0.5721 0.5815 0.5768
40% 0.8683 0.5618 0.6822 0.4683 0.4735 0.4709
100% 0.8565 0.4152 0.5593 0.4041 0.4056 0.4048
Table 11: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model with CoT. We perform limited ablations because of the insights from the non CoT model training.
Fraction of NV data Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
2% 0.9226 0.8901 0.906 0.7688 0.7638 0.7663
5% 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Table 12: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Llama3.2-3B-Instruct model. We perform limited ablations because of the insights from the non CoT model training.
Fraction of NV data Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
1% 0.654 0.6468 0.6504 0.491 0.4788 0.4848
2% 0.6636 0.6057 0.6333 0.4869 0.4745 0.4806
Table 13: Effect of Simple Ensembling Strategies on Detection and Localization Performance: Comparison of intersection and union-based ensembling of SFT and RS-SFT models on Qwen3-4B. The intersection strategy retains only line numbers predicted by both models, while the union strategy aggregates all predicted line numbers from either model. Both ensembling approaches underperform the individual models, with the intersection strategy exhibiting particularly poor precision and recall, suggesting that the observed precision–recall differences are driven by training setup rather than a fundamental trade-off.
Ensembling Strategy Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B DPO Intersection (non-CoT & CoT) 0.0667 0.0033 0.0063 0.0000 0.0000 0.0000
Qwen3-4B DPO Union (non-CoT & CoT) 0.6188 0.4187 0.4995 0.1952 0.1251 0.1524
Table 14: Effect of Removing NO VIOLATIONS (NV) Data on Hard PEP Benchmark Performance: Full detection and localization results for Qwen3-4B trained with SFT+RS-DPO under different NV data settings. The 0% NV variant achieves substantially higher recall than the model reported in the main paper, with only a moderate reduction in precision. Since the precision drop is smaller than the recall gain, this leads to an improved overall F score, confirming that the originally reported recall was an underestimate and further strengthening our state-of-the-art recall claim.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B (SFT + RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B (SFT + RS-DPO, 0% NV) 0.6658 0.8037 0.7283 0.3797 0.2063 0.2674
Qwen3-4B w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
F.4 PEP Benchmark Creation Additional Details
As discussed in section 4.3 we use some high recall heuristics to find promising candidates for detecting the selected hard PEP best practice idioms. These are summarized in Table 15, 16 and 17.
Table 15: High recall heuristics used to find instances of PEP violations that human annotators vet
PEP Description Heuristics Example
506
Adds secrets module
to the standard library
for cryptographically secure
random value generation
Conjunction of 2 conditions:
1. Presence of "random" module imports
2. Presence of "random" function usage
characters = string.ascii_letters +
string.punctuation + string.digits
password = "".join(random.choice
(characters) for x in range(16))
Use instead:
characters = string.ascii_letters +
string.punctuation + string.digits
password = "".join(secrets.choice
(characters) for x in range(16))
557
Introduces the dataclasses
module, enabling automatic
generation of common boilerplate
methods for classes
Conjunction of 2 conditions:
1. There is a class with manual
implementation of "__init__" method
2. On the same class there is manual
implementation of common
special methods or comparison methods
that follow standard data storage patterns.
"class Point:
def __init__(self, x, y):
self.x = x
self.y = y
def __repr__(self):
return f""Point(x={self.x}, y={self.y})""
Use instead:
from dataclasses import dataclass
@dataclass
class Point:
x: int
y: int"
655
Introduces Required[] and
NotRequired[] type qualifiers
to replaces cumbersome
TypedDict inheritance patterns.
Conjuction of:
1. "TypedDict" defined with inheritance pattern.
2. total=False parameter usage in class definition
class _MovieBase(TypedDict): # implicitly total=True
title: str
class Movie(_MovieBase, total=False):
year: int
Use instead:
class Movie(TypedDict):
title: str
year: NotRequired[int]
634
Introduced structural pattern
matching, enabling more
expressive and concise ways
to match data structures
and control flow.
Multiple consecutive if-elif-else
statements that compare a
single variable against different
values with dysjunction of 2 conditions:
1. Length of ladder (number of
conditons at the "top level" + one level in) >= 6
2. Depth of ladder (degree of nesting) >=3
"def handle_response(response):
if isinstance(response, dict):
if ""error"" in response:
print(f""Error: {response['error']}"")
elif ""data"" in response:
print(f""Data: {response['data']}"")
else:
print(""Unknown response format"")
elif isinstance(response, list):
print(""List of items:"", response)
else:
print(""Invalid response type"")
Use instead:
def handle_response(response):
match response:
case {""error"": error_message}:
print(f""Error: {error_message}"")
case {""data"": data_content}:
print(f""Data: {data_content}"")
case list(items):
print(""List of items:"", items)
case _: print(""Invalid response type"")"
614
Removes previous restrictions
on decorator syntax. Before,
only simple names or dotted
names were valid decorators.
After 614, any valid expression
can be used as a decorator
Conjunction of 2 conditions:
1. A decorator is applied using a name
(e.g., @decorator) where that name is
assigned earlier in the code.
2. The assignment value is an expression
of type Call, Attribute, or Subscript (e.g.,
deco = factory(), deco = module.decorator,
deco = decorators[i]).
# def uppercase(func):
def wrapper(*args, **kwargs):
return func(*args, **kwargs).upper()
return wrapper
@uppercase
def greet():
return "hello"
Use Instead:
deco = [uppercase]
@deco[0]
def greet2():
return "hi"
616
Replaces manual slicing
with dedicated methods
dysjunction of 2 conditions:
1. There is a "check" with
startswith or endswith on a
given variable x.
2. On the same variable x
check if there is an "edit" using a
program slicing syntax or using "replace()".
if s.startswith(prefix): s = s[len(prefix):]
Use instead:
s = s.removeprefix(prefix)
OR
s[:-len(suffix)]
Use instead:
s.removesuffix(suffix)
584
Introduces the binary operators
| (merge) and |= (update) on
dict (and other built-in mapping
types), providing an expressive,
in-place-or-new‐object way to
combine dictionaries.
disjunction of two conditions:
1. A copy-and-update sequence
on the same variable
or in close proximity: d = d1.copy()
followed by d.update(d2)
2. A dictionary literal using multiple
unpackings {**d1, **d2},
indicating ad-hoc merging rather
than the new operators
d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
merged = d1.copy()
merged.update(d2)
d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
merged = {**d1, **d2}
d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
merged = dict(list(d1.items()) + list(d2.items()))
Use instead:
d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
merged = d1 | d2
d1 = {'a': 1, 'b': 2} d2 = {'c': 3, 'd': 4}
d1 |= d2 # d1 is now {'a': 1, 'b': 2, 'c': 3, 'd': 4}
Table 16: High recall heuristics used to find instances of PEP violations that human annotators vet
PEP Description Heuristics Example
570
Introduces new syntax
(the / marker) in Python
function signatures to
specify positional-only
parameters, ensuring that
certain arguments can
only be supplied by their
position and not as keywords
Conjunction of the
following conditions:
1. Have only positional-or-
keyword parameters
(without *args, **kwargs, keyword
-only parameters, or the '/' marker),
2. Include 2 to 4 parameters, all of
which have no default values
def compute_area(width, height):
return width * height
area = compute_area(width=5, height=10)
print("Area:", area)
Use instead:
def compute_area(width, height, /):
return width * height
area = compute_area(5, 10)
print("Area:", area)
567
Adds the contextvars module,
enabling context-local variables
for managing dynamic state.
Dysjunction of the following conditions:
1. Look for import threading together
with threading.local() object creation and use.
2. Find global statements or assignment
to variables at the module
level that are accessed or mutated in functions,
especially as shared state.
3. Identify async functions or classes
where context or state
variables are passed as parameters
(e.g., def func(context, …) or async
def func(context, …)), not as context-local variables.
import threading
_thread_local = threading.local()
def set_context(value):
_thread_local.value = value
def get_context():
return getattr(_thread_local, 'value', None)
Use instead:
from contextvars import ContextVar
context_var = ContextVar('value')
def set_context(value):
context_var.set(value)
def get_context():
return context_var.get()
530
Enables the use of "async for"
and "await" in list, set, and dict
comprehensions as well as in
generator expressions, providing
concise asynchronous data
processing within comprehensions
Dysjunction of the following conditions:
1. "async" def functions that uses "async for"
loops to build lists, sets, or dicts.
2. "async for" loops, followed by methods like
result.append(…), result.extend(…), or result[key] = ….
3. Comprehensions written without the "async for"
clause despite being inside an "async def"
result = []
async for i in aiter():
if i % 2:
result.append(i)
Use instead:
result = [i async for i in aiter() if i % 2]
525
Introduces the ability to define
asynchronous generator functions
using the async def and yield
syntax, enabling concise, native
support for asynchronous iteration.
Dysjunction of the following conditions:
1. classes defining both "__aiter__" and "__anext__"
methods, especially where the class is used solely to
produce a sequence of values asynchronously.
2. async def functions that create and return custom
iterator classes instead of using async def with yield.
class Ticker:
"""Yield numbers from 0 to `to`
every `delay` seconds."""
def __init__(self, delay, to):
self.delay = delay
self.i = 0
self.to = to
def __aiter__(self):
return self
async def __anext__(self):
i = self.i
if i >= self.to:
raise StopAsyncIteration
self.i += 1
if i:
await asyncio.sleep(self.delay)
return i
Use instead:
async def ticker(delay, to):
"""Yield numbers from 0 to `to`
every `delay` seconds."""
for i in range(to):
yield i
await asyncio.sleep(delay)
520
Ensures that the order in
which attributes are defined within
a class body is preserved in the
resulting class object, making
the attribute order predictable
and consistent.
Dysjunction of the following conditions:
1. Uses sorted() or otherwise processes
class.__dict__.keys() to impose attribute order.
2. Attribute names are tracked in a list or similar
structure solely to maintain definition order.
3. Custom metaclass logic or "__prepare__" impleme-
ntations created to preserve the order of class attributes.
class Person:
name = "Alice"
age = 30
city = "Wonderland"
def display_attributes(self):
# Manually sorting keys
for key in sorted(self.__class__.__dict__.keys()):
if not key.startswith("__"):
print(key, getattr(self, key))
Use instead:
class Person:
name = "Alice"
age = 30
city = "Wonderland"
def display_attributes(self):
# Directly iterate over the preserved definition order
for key in self.__class__.__definition_order__:
print(key, getattr(self, key))
498
Introduces f-strings (formatted
string literals) as a new,
concise, and efficient way
to embed Python
expressions inside string
literals using the f'' prefix.
Dysjunction of the following conditions:
1. Occurrences of string literals with
.format(…) applied,
especially where keys or variables match
braces in the string
2. String literals concatenated using
"+" with variables.
3. Uses of the "%" operator for string formatting,
name = "Alice"
age = 30
greeting = "Hello, " + name +
"! You are " + str(age) + " years old."
Use instead:
name = "Alice"
age = 30
greeting = f"Hello, {name}!
You are {age} years old."
OR
value = 12.3456
formatted = "The value is {:.2f}".format(value)
Use instead:
value = 12.3456
formatted = f"The value is {value:.2f}"
Table 17: High recall heuristics used to find instances of PEP violations that human annotators vet
PEP Description Heuristics Example
487
Makes customizing class
creation and subclass initialization
easier by introducing __init_
subclass__ and __set_name__,
eliminating the need for
most custom metaclasses
Disjunction of the following conditions:
1. Custom metaclasses defined to
execute code during class creation
or subclassing (e.g., overriding __new__,
__init__, or __call__ in metaclasses)
instead of using __init_subclass__.
2. Descriptor classes lacking __set_
name__ method and employing
manual workarounds to determine
their assigned attribute names.
3. Classes or frameworks manually
tracking or registering sub-
classes via metaclass hooks instead
of leveraging __init_subclass__.
class Meta(type):
def __new__(meta, name, bases, namespace):
for key, value in namespace.items():
if isinstance(value, Descriptor):
value.name = key
return super().__new__
(meta, name, bases, namespace)
class MyClass(metaclass=Meta):
attr = Descriptor()
Use instead:
class Descriptor:
def __set_name__(self, owner, name):
self.name = name
class MyClass:
attr = Descriptor()
OR
class PluginBase(type):
plugins = {}
def __new__(meta, name, bases, namespace):
if name != 'Plugin':
meta.plugins[name] = namespace['priority']
return super().__new__
(meta, name, bases, namespace)
class Plugin(metaclass=PluginBase):
priority = 0
class HighPriority(Plugin):
priority = 10
Use instead:
class Plugin:
plugins = {}
priority = 0
def __init_subclass__(cls, **kwargs):
super().__init_subclass__(**kwargs)
cls.plugins[cls.__name__] = cls.priority
class HighPriority(Plugin):
priority = 10
593
Introduces flexible function
and variable annotations via
typing.Annotated, which lets you
attach context-specific metadata
to type hints (e.g., validation
constraints, units)
Conjunction of 2 conditions:
1. Type hints are already present
in function arguments, return
types, or variable annotations.
2. Nearby comments/docstrings (within
±2 lines) contain metadata-like
patterns such as "min", "max",
"nullable", "regex", "enum",
"unit", "deprecated", etc.
# max 100, min 1
def set_age(age: int) ->None:
pass
Use instead:
from typing import Annotated
Age = Annotated[int, "min=1", "max=100"]
def set_age(age: Age) ->None:
pass
526
introduces explicit variable
annotations, allowing type
hints directly on variable
declarations for local, global,
and class variables in Python
Disjunction of the following conditions:
1. Variables assigned values with a
type comment (e.g., x = 0 # type: int)
instead of using annotation syntax.
2. Identify variable assignments,
especially class and
instance attributes, that lack any
type annotation (e.g., name = "" in class bodies).
3.Module-level variables assigned
values without accompanying type hints—
especially in type-annotated codebases.
# type: List[int]
numbers = []
Use instead:
numbers: List[int] = []
OR
class Player:
# type: str<br>name = "Guest"
Use instead:
class Player:
name: str = "Guest"
589
Introduces TypedDict, enabling
precise type hints for
dictionaries with a fixed
set of string keys, improving
static type checking and
readability in Python code.
Disjunction of the following conditions:
1. Dictionary literals or variables
consistently using the same fixed set of
string keys without accompanying
TypedDict annotations.
2. Functions annotated with broad
dictionary types like Dict[str, Any], dict,
or untyped parameters/returns that actually
expect dictionaries with a known fixed set of keys.
3. Explicit key presence checks or
accessing dictionary keys repeatedly that suggest
a structured dictionary shape.
movie = {'name': 'Blade Runner',
'year': 1982}
Use instead:
from typing import TypedDict
class Movie(TypedDict):
name: str
year: int
movie: Movie = {'name': 'Blade Runner',
'year': 1982}
572
Introduces the assignment
expression operator :=
(the "walrus operator"),
allowing assignment to
variables within expressions,
1. Patterns where a value is first assigned
to a variable, and then immediately checked
or used in the next line or inside a
loop, list comprehension, or condition.
2. separate assignment and conditional test statements
match = pattern.search(data)
if match is not None:
process(match)
Use instead:
if (match := pattern.search(data)) is not None:
process(match)
Appendix G More Results
Table 18: Cross-Idiom Generalization on Python Ruff Best Practice Idioms by Transfer Setting: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff best practice idioms, and the performance is reported on other Ruff best practice idioms with varying levels of transfer - In-Domain, Near Transfer, and Far Transfer (section 4.2).
Model In-Domain Near Transfer Far Transfer
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Qwen3-4B 0.45 0.14 0.22 0.58 0.24 0.34 0.54 0.29 0.38
+SFT 0.93 (+0.48) 0.74 (+0.6) 0.83 (+0.61) 0.89 (+0.31) 0.24 (+0) 0.38 (+0.04) 0.72 (+0.18) 0.27 (-0.02) 0.39 (+0.01)
+RS-DPO 0.72 (+0.27) 1 (+0.86) 0.83 (+0.61) 0.76 (+0.18) 0.8 (+0.56) 0.78 (+0.44) 0.75 (+0.21) 0.81 (+0.52) 0.78 (+0.4)
Qwen3-4B w CoT 0.87 0.5 0.63 0.95 0.88 0.91 0.87 0.68 0.76
+RS-SFT 0.87 (+0) 0.73 (+0.23) 0.8 (+0.17) 0.97 (+0.02) 0.86 (-0.02) 0.91 (+0) 0.94 (+0.07) 0.82 (+0.14) 0.88 (+0.12)
+RS-DPO 0.86 (-0.1) 0.85 (+0.35) 0.85 (+0.22) 0.97 (+0.02) 0.92 (+0.04) 0.94 (+0.03) 0.92 (+0.05) 0.86 (+0.18) 0.89 (+0.13)
Llama3.2-3B-Instruct 0.54 0.43 0.48 0.69 0.68 0.69 0.47 0.51 0.49
+SFT 0.88 (+0.34) 0.87 (+0.44) 0.88 (+0.4) 0.89 (+0.2) 0.44 (-0.24) 0.59 (-0.1) 0.61 (+0.14) 0.27 (-0.24) 0.37 (-0.12)
+RS-DPO 0.75 (+0.21) 0.92 (+0.49) 0.83 (+0.35) 0.81 (+0.12) 0.71 (+0.03) 0.76 (+0.07) 0.61 (+0.14) 0.59 (+0.08) 0.60 (+0.11)
Table 19: Ablation on In-Context Examples for the Hard PEP Benchmark: Effect of including before/after code examples in the prompt for base Qwen3-4B models and final MetaLint models trained with SFT and DPO. Across most settings, removing examples results in minimal changes in detection and localization performance, indicating that the natural-language descriptions alone are often sufficient. The only notable degradation occurs for the CoT-trained MetaLint model, which exhibits a modest drop in detection performance without examples due to an increased tendency to predict NO VIOLATIONS, suggesting that this more conservative variant benefits more from additional in-context guidance.
Model Examples? Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B Yes 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B w CoT Yes 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-4B (SFT + RS-DPO) Yes 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B w CoT (RS-SFT + RS-DPO) Yes 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Qwen3-4B No 0.5556 0.1736 0.2645 0.0802 0.0733 0.0766
Qwen3-4B w CoT No 0.8304 0.3669 0.5089 0.2572 0.1542 0.1928
Qwen3-4B (SFT + RS-DPO) No 0.7167 0.7268 0.7217 0.3396 0.1909 0.2444
Qwen3-4B w CoT (RS-SFT + RS-DPO) No 0.8215 0.4249 0.5601 0.2860 0.1983 0.2342
G.1 Expanded Results on the PEP Hard Best Practice Benchmark
We show the expanded results across various model sizes for the evaluated model families in Table 20. We note that most results follow the expected trends with more parameters or CoT usage leading to better performance but there are soem exceptions to the trend. We mainly see this for cases like Qwen2.5 and Qwen2.5Coder families. We note that Qwen2.5Coder-7B-Instruct has almost zero metrics because it always predicts NO VIOLATIONS FOUND for all instances and Qwen2.5Coder-14B-Instruct has really low scores because of similar reasons. for Qwen2.5 family we notice that 32B variant performs a bit worse than 32B.
We also analyze MetaLint SFT models on the hard PEP benchmark and observe that they perform similarly or slightly worse than the base untrained models. This suggests that SFT alone may lead to overfitting on the Ruff best practice idiom distribution and struggles to generalize from easy to hard cases without DPO training. These findings highlight the importance of the DPO (preference-tuning) stage in the MetaLint pipeline. However, we also emphasize that while the SFT stage can limit generalization, it remains essential for effective DPO training, as it teaches the LLM to follow the correct output format and establishes a strong base policy. This is supported by our experiments with the CoT model, where applying RS-DPO directly to the Qwen/Qwen3-4B model (without SFT) led to near-zero performance across all metrics, as the model consistently failed to produce outputs in the required format.
Table 20: Results on the hard PEP benchmark to measure easy to hard generalization.
Detection Localization
Model $P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Llama3.2-3B-Instruct 0.7042 0.214 0.3283 0.0691 0.0798 0.0741
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B with CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-8B 0.8267 0.3572 0.4988 0.1806 0.1285 0.1501
Qwen3-8B with CoT 0.8886 0.4672 0.6124 0.3122 0.2029 0.2459
Qwen3-14B 0.9021 0.4612 0.6103 0.289 0.2521 0.2693
Qwen3-14B with CoT 0.9116 0.4857 0.6337 0.3993 0.2915 0.3369
Qwen3-32B 0.9021 0.5205 0.6601 0.2807 0.2711 0.2758
Qwen3-32B with CoT 0.9377 0.5645 0.7048 0.4152 0.3086 0.354
Qwen2.5-3B-Instruct 0.0667 0.0033 0.0063 0.0036 0.0036 0.0036
Qwen2.5-7B-Instruct 0.4333 0.1379 0.2092 0.0585 0.0518 0.0549
Qwen2.5-14B-Instruct 0.8017 0.4324 0.5618 0.2389 0.2158 0.2267
Qwen2.5-32B-Instruct 0.8667 0.2656 0.4066 0.163 0.1477 0.155
Qwen2.5Coder-3B-Instruct 0.7802 0.411 0.5384 0.1257 0.0745 0.0936
Qwen2.5Coder-7B-Instruct 0.0667 0.0033 0.0063 0 0 0
Qwen2.5Coder-14B-Instruct 0.2 0.0443 0.0726 0.0294 0.0264 0.0278
Qwen2.5Coder-32B-Instruct 0.8961 0.5328 0.6683 0.3432 0.3077 0.3245
DeepSeek-R1-Distill-Qwen-7B with CoT 0.7143 0.2841 0.4065 0.1064 0.1122 0.1092
DeepSeek-R1-Distill-Qwen-14B with CoT 0.69 0.2345 0.35 0.1856 0.1245 0.149
DeepSeek-R1-Distill-Qwen-32B with CoT 0.9008 0.5899 0.713 0.4015 0.3403 0.3684
GPT-oss-20b 0.8377 0.3531 0.4968 0.251 0.1695 0.2024
GPT-oss-120b 0.9157 0.6456 0.7573 0.3991 0.3331 0.3631
Qwen3-4B MetaLint (SFT) (Ours) 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
Qwen3-4B MetaLint (SFT+RS-DPO) (Ours) 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
Qwen3-4B MetaLint w CoT (RS-SFT) (Ours) 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
Qwen3-4B MetaLint w CoT (RS-SFT+RS-DPO) (Ours) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Llama3.2-3B-Instruct MetaLint (SFT) (Ours) 0.5627 0.259 0.3547 0.1066 0.0509 0.0689
Llama3.2-3B-Instruct MetaLint (SFT+RS-DPO) (Ours) 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
o3-mini 0.8939 0.5845 0.7068 0.3169 0.2361 0.2706
o4-mini 0.9667 0.5943 0.7361 0.4131 0.3164 0.3584
GPT-4o 0.8938 0.6788 0.7716 0.4461 0.332 0.3807
GPT-4.1 0.907 0.646 0.7546 0.4632 0.4673 0.4653
GPT-5 (high) 0.913 0.5673 0.6998 0.4397 0.4257 0.4326
Table 21: Cross-Idiom Generalization on Python Ruff Best Practice Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff best practice idioms and tested on other Ruff best practice idioms with varying levels of transfer (section 4.2). Best score across the compared training setups per model is bolded.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.538 0.2637 0.3539 0.1396 0.1479 0.1436
Qwen3-4B + SFT 0.7686 0.3178 0.4497 0.2976 0.296 0.2968
Qwen3-4B + SFT + RS-DPO 0.7469 0.8315 0.7869 0.6527 0.6696 0.6611
Qwen3-4B w CoT 0.8812 0.6854 0.771 0.5049 0.4878 0.4962
Qwen3-4B w CoT + RS-SFT 0.935 0.8183 0.8727 0.6639 0.65 0.6569
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9234 0.8643 0.8929 0.771 0.7571 0.764
Llama3.2-3B-Instruct 0.5092 0.5286 0.5187 0.1371 0.3 0.1882
Llama3.2-3B-Instruct + SFT 0.6793 0.3598 0.4704 0.3424 0.3485 0.3454
Llama3.2-3B-Instruct + SFT + RS-DPO 0.654 0.6468 0.6504 0.491 0.4788 0.4848
Table 22: Easy-to-Hard Generalization on PEP Best Practice Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff best practice idioms and tested on hard manually curated PEP best practice violation detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B 0.5267 0.1715 0.2587 0.0954 0.0824 0.0884
Qwen3-4B + SFT 0.4333 0.0821 0.1381 0.0432 0.0221 0.0292
Qwen3-4B + SFT + RS-DPO 0.7031 0.7043 0.7037 0.3536 0.193 0.2497
Qwen3-4B w CoT 0.8154 0.3986 0.5354 0.2625 0.1467 0.1882
Qwen3-4B w CoT + RS-SFT 0.7615 0.3689 0.497 0.2785 0.1437 0.1896
Qwen3-4B w CoT + RS-SFT + RS-DPO 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
Llama3.2-3B-Instruct 0.7042 0.214 0.3283 0.0691 0.0798 0.0741
Llama3.2-3B-Instruct + SFT 0.5627 0.259 0.3547 0.1066 0.0509 0.0689
Llama3.2-3B-Instruct + SFT + RS-DPO 0.6368 0.5614 0.5965 0.2364 0.1263 0.1647
G.2 Statistical Significance of Results on the PEP Hard Best Practice Benchmark
To analyze the statistical significance of performance differences over the PEP benchmark, we conduct Wilcoxon signed-rank tests comparing various MetaLint variants against each other and against baseline models. We evaluate instance-level detection accuracy (binary labels indicating whether the LLM correctly predicted the presence of a violation) as well as instance-level precision and recall for line-level localization. To control for multiple comparisons, we apply a Bonferroni correction to adjust the significance threshold $\alpha$ as $\alpha=\frac{0.05}{m}$ where $m$ is the number of comparisons (or rows in any given statistical significance table in this case).
Table 24 reports the Wilcoxon signed-rank test statistic and corresponding $p$-value (in parentheses) for detection accuracy, localization precision, and localization recall when comparing various MetaLint variants to assess the effects of RS-DPO and CoT. We find that applying RS-DPO to the base SFT policy leads to statistically significant improvements in both detection and localization performance, with RS-DPO consistently outperforming the original SFT checkpoint across all three metrics with it being always better for localization. For the CoT variant, RS-DPO also yields consistent but less significant gains, likely because the RS-SFT CoT checkpoint is already relatively strong. Finally, we observe no statistically significant difference between the CoT (RS-SFT+RS-DPO) and the standard (SFT+RS-DPO) variant, suggesting that CoT does not provide a meaningful additional benefit in this setting.
Table 25 shows the statistical significance of comparing the base untrained model Qwen3-4B with its MetaLint variants (SFT and SFT+RS-DPO), and the Qwen3-4B CoT model with MetaLint w/ CoT (RS-SFT and RS-SFT+RS-DPO). The SFT variant yields significant gains in detection and localization recall, but not in localization precision. The SFT+RS-DPO model improves significantly across all three metrics. In contrast, training RS-SFT from the Qwen3-4B w/ CoT base does not yield significant improvements. However, the RS-SFT+RS-DPO variant produces significant gains in localization precision and recall, but not detection. These results suggest that while SFT alone offers limited generalization, combining it with DPO reliably improves localization and can significantly boost detection when starting from a weaker base model.
Table 26 shows the statistical significance results when comparing the MetaLint (SFT+RS-DPO) and MetaLint w CoT (RS-SFT+RS-DPO) variants against various baselines. Here we want to higlight that MetaLint offers comparable performance across two out of three or all three metrics against several 32B models that outperform it like Qwen3-32B, Qwen3-32B w CoT, Qwen2.5Coder-32B and R1-Distill-Qwen-32B. Also the MetaLint non CoT (SFT+RS-DPO) variant has no singificant difference in performance compared to o3-mini, soldifying that MetaLint without CoT has generalized to the point of being as capable as o3-mini (even though the Qwen3-4B models without CoT and Qwen3-4B model with CoT perform worse than it with the difference being statistically singificant in Table 23).
Table 23: Wilcoxon signed-rank test results comparing untrained Qwen3-4B variants with o3-mini, using Bonferroni-adjusted significance threshold $\alpha=0.025$. Each cell reports the test statistic (p-value).
Model Comparison Detection Localization P Localization R
Qwen3-4B vs o3-mini 1266.5 (7.20e-21) 743.5 (2.96e-11) 739.0 (4.23e-09)
Qwen3-4B w CoT vs o3-mini 921.5 (3.23e-09) 2385.5 (9.61e-02) 1891.0 (4.99e-04)
Table 27 shows the effect of using a CoT for the Qwen3 model families and we notice that using a CoT leads to singificant gains for all metrics for the 4B and 8B models indicating that for smaller models CoTs might be essential for good performance on this task. However the 14B and 32B model only show statistically significant improvement in localization precision with the CoT indicating that the CoT might offer limited benefit for larger models.
Table 28 shows the effect of varying model scale for the Qwen3, Qwen2.5, Qwen2.5Coder, and DeepSeek-R1-Distill-Qwen families. For Qwen3 we see benefits moving from 4B to 8B abd 8B to 14B but no statistically significant difference moving from 14B to 32B when not using a CoT. Wehn using a CoT for Qwen3 we notice that the performance differences are rarely different in terms of statistical significant except for localizaiton performance between 4B and 8B and 8B and 14B. For R1-Distill-Qwen family we notice a significant difference moving from 14B to 32B but not for 7B to 14B. For the Qwen2.5Coder family we notice difference across all model scales, but the trend is weird with a big drop in performance from 3B to 7B and then a slow climb back to great performance around 32B. We notice that for the Qwen2.5 family which shows relatively reasonable trends with model scale, the performance differences are statistically singificant execpt for the performance gain from 14B to 32B being significant only for recall. To conclude the trends across model scales vary a lot across model families but in general the model size does help but differences may be smaller if the models are capable of reasoning and use a CoT.
Table 29 shows comparison between the GPT models. We only compared GPT-4o and its successor GPT-4.1 and o3-mini against o4-mini and the results show that GPT-4.1 is only significantly better for localization recall while o4-mini is beter than o3-mini for overall localization but not for detection.
Table 24: Wilcoxon signed-rank test results comparing MetaLint variants. Each cell reports test statistic (p-value). All the MetaLint models are trained Qwen3-4B variants. We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
Model Comparison Detection Localization P Localization R
MetaLint (SFT) vs MetaLint (SFT+RS-DPO) 7192.0 (1.92e-12) 0.0 (2.49e-20) 0.0 (2.92e-18)
MetaLint w CoT (RS-SFT) vs MetaLint w CoT (RS-SFT+RS-DPO) 839.5 (2.18e-03) 740.0 (3.95e-03) 523.0 (2.34e-05)
MetaLint (SFT) vs MetaLint w CoT (RS-SFT) 528.0 (6.91e-14) 11.0 (1.38e-15) 113.0 (7.95e-12)
MetaLint (SFT+RS-DPO) vs MetaLint w CoT (RS-SFT+RS-DPO) 8140.0 (5.55e-01) 2568.5 (8.42e-01) 2544.0 (4.44e-01)
Table 25: Wilcoxon signed-rank test results comparing MetaLint models against their untrained counterparts, with Bonferroni-adjusted significance threshold $\alpha=0.0125$. Each cell reports the test statistic (p-value).
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-4B MetaLint (SFT) 560.5 (8.64e-03) 363.5 (1.82e-02) 238.0 (4.85e-04)
Qwen3-4B vs Qwen3-4B MetaLint (SFT+RS-DPO) 7260.0 (4.13e-09) 411.0 (1.99e-15) 979.5 (7.30e-09)
Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT) 1224.0 (7.22e-01) 937.0 (6.12e-01) 918.5 (5.39e-01)
Qwen3-4B w CoT vs Qwen3-4B MetaLint w CoT (RS-SFT+RS-DPO) 1728.0 (1.83e-02) 1011.0 (1.53e-03) 966.0 (1.02e-03)
Table 26: Wilcoxon signed-rank test statistics and p-values comparing MetaLint variants against baseline models. All the MetaLint variants are Qwen3-4B variants and Qwen2.5 and Qwen2.5Coder variants are instruction tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0017$.
Model Comparison Detection Localization P Localization R
Qwen3-8B vs MetaLint (SFT+RS-DPO) 8140.5 (7.21e-03) 1309.5 (2.19e-08) 2067.0 (2.12e-03)
Qwen3-8B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2070.0 (9.17e-01) 1974.5 (1.88e-01) 2161.0 (6.58e-01)
Qwen3-14B vs MetaLint (SFT+RS-DPO) 7304.0 (4.96e-01) 2816.5 (3.15e-02) 3159.0 (2.88e-02)
Qwen3-14B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2392.0 (2.78e-01) 2749.5 (1.26e-01) 2319.0 (1.32e-03)
Qwen3-32B vs MetaLint (SFT+RS-DPO) 7175.0 (4.48e-01) 3262.0 (1.93e-02) 2818.5 (5.56e-03)
Qwen3-32B w CoT vs MetaLint w CoT (RS-SFT+RS-DPO) 2677.5 (9.95e-03) 3479.0 (8.38e-02) 3180.5 (5.64e-04)
R1-Distill-Qwen-7B vs MetaLint (SFT+RS-DPO) 8244.0 (2.65e-08) 555.0 (8.12e-15) 1924.0 (1.91e-04)
R1-Distill-Qwen-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 2907.0 (7.07e-10) 915.5 (1.04e-12) 1569.5 (8.64e-06)
R1-Distill-Qwen-14B vs MetaLint (SFT+RS-DPO) 9877.0 (3.99e-06) 2582.5 (9.36e-06) 3085.0 (2.00e-03)
R1-Distill-Qwen-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 2660.0 (9.11e-08) 1703.0 (2.98e-06) 1791.0 (3.97e-05)
R1-Distill-Qwen-32B vs MetaLint (SFT+RS-DPO) 8677.5 (2.51e-01) 5767.5 (1.93e-01) 3705.0 (6.67e-06)
R1-Distill-Qwen-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 3125.0 (3.11e-02) 3641.5 (6.35e-02) 2175.0 (4.99e-06)
Qwen2.5-3B vs MetaLint (SFT+RS-DPO) 8001.0 (1.41e-15) 0.0 (1.24e-22) 68.5 (1.26e-19)
Qwen2.5-3B vs MetaLint w CoT (RS-SFT+RS-DPO) 949.0 (4.96e-23) 0.0 (4.24e-22) 0.0 (9.89e-20)
Qwen2.5-7B vs MetaLint (SFT+RS-DPO) 7312.5 (3.37e-10) 208.0 (1.44e-18) 610.0 (1.99e-12)
Qwen2.5-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 1187.5 (1.14e-14) 226.0 (2.60e-18) 406.5 (5.69e-14)
Qwen2.5-14B vs MetaLint (SFT+RS-DPO) 8677.5 (2.51e-01) 3045.5 (5.70e-04) 4006.0 (3.83e-01)
Qwen2.5-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 4123.0 (4.86e-01) 3228.0 (1.51e-03) 4383.0 (9.89e-01)
Qwen2.5-32B vs MetaLint (SFT+RS-DPO) 8640.0 (1.76e-04) 1492.5 (3.12e-08) 2971.5 (4.55e-02)
Qwen2.5-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 1792.0 (8.16e-06) 1166.0 (2.43e-08) 1983.5 (4.71e-03)
Qwen2.5Coder-3B vs MetaLint (SFT+RS-DPO) 11184.0 (8.64e-03) 953.5 (3.73e-12) 1716.0 (3.00e-07)
Qwen2.5Coder-3B vs MetaLint w CoT (RS-SFT+RS-DPO) 3683.5 (6.45e-03) 1126.0 (4.39e-11) 1403.5 (1.32e-08)
Qwen2.5Coder-7B vs MetaLint (SFT+RS-DPO) 8001.0 (1.41e-15) 0.0 (7.69e-23) 0.0 (2.02e-20)
Qwen2.5Coder-7B vs MetaLint w CoT (RS-SFT+RS-DPO) 949.0 (4.96e-23) 0.0 (2.61e-22) 0.0 (6.55e-20)
Qwen2.5Coder-14B vs MetaLint (SFT+RS-DPO) 9123.5 (1.04e-12) 289.0 (8.19e-20) 736.0 (7.32e-14)
Qwen2.5Coder-14B vs MetaLint w CoT (RS-SFT+RS-DPO) 1112.0 (1.82e-19) 159.5 (7.63e-20) 408.5 (3.02e-15)
Qwen2.5Coder-32B vs MetaLint (SFT+RS-DPO) 6833.5 (2.86e-01) 4500.0 (9.59e-01) 2651.5 (7.07e-05)
Qwen2.5Coder-32B vs MetaLint w CoT (RS-SFT+RS-DPO) 1039.5 (1.16e-02) 2655.0 (9.39e-01) 1235.5 (1.44e-05)
o3-mini vs MetaLint (SFT+RS-DPO) 7520.0 (4.83e-02) 4986.0 (5.20e-01) 4427.5 (2.87e-01)
o3-mini vs MetaLint w CoT (RS-SFT+RS-DPO) 1944.0 (7.15e-04) 3169.0 (5.16e-01) 2683.0 (4.23e-01)
Table 27: Wilcoxon signed-rank test results measuring the effect of Chain-of-Thought (CoT) prompting across Qwen3 model scales. Each cell reports test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.0125$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-4B w CoT 1260.0 (3.99e-08) 425.0 (1.91e-09) 624.5 (2.25e-05)
Qwen3-8B vs Qwen3-8B w CoT 1924.0 (4.27e-03) 1132.5 (8.69e-06) 1005.0 (1.15e-04)
Qwen3-14B vs Qwen3-14B w CoT 1691.0 (2.01e-01) 2127.0 (4.27e-04) 2398.5 (1.26e-01)
Qwen3-32B vs Qwen3-32B w CoT 1572.5 (2.75e-01) 1767.0 (6.35e-06) 2596.5 (7.31e-02)
Table 28: Wilcoxon signed-rank test results measuring the effect of increasing model scale across families and CoT settings. Each cell shows the test statistic (p-value). All Qwen2.5 and Qwen2.5Coder variants are instruction-tuned checkpoints. We use the Bonferroni corrected significance threshold $\alpha=0.0036$.
Model Comparison Detection Localization P Localization R
Qwen3-4B vs Qwen3-8B 546.0 (2.35e-08) 618.0 (1.37e-04) 572.5 (1.72e-03)
Qwen3-8B vs Qwen3-14B 1350.0 (2.11e-03) 1503.5 (3.96e-05) 1008.5 (2.69e-07)
Qwen3-14B vs Qwen3-32B 875.0 (2.22e-02) 3129.5 (7.94e-01) 2061.0 (4.14e-01)
Qwen3-4B w CoT vs Qwen3-8B w CoT 1468.5 (1.90e-02) 1578.5 (1.07e-01) 1081.5 (2.60e-03)
Qwen3-8B w CoT vs Qwen3-14B w CoT 904.5 (1.40e-01) 1248.0 (1.66e-03) 1099.5 (1.82e-05)
Qwen3-14B w CoT vs Qwen3-32B w CoT 850.0 (3.78e-02) 1834.5 (4.87e-01) 2396.0 (4.40e-01)
R1-Distill-Qwen-7B vs R1-Distill-Qwen-14B 4278.0 (2.67e-01) 1431.0 (4.66e-03) 1962.5 (6.01e-01)
R1-Distill-Qwen-14B vs R1-Distill-Qwen-32B 2432.0 (1.44e-12) 1475.5 (8.07e-11) 843.5 (6.38e-15)
Qwen2.5Coder-3B vs Qwen2.5Coder-7B 1541.0 (4.56e-14) 0.0 (3.46e-10) 0.0 (7.07e-10)
Qwen2.5Coder-7B vs Qwen2.5Coder-14B 8.0 (7.89e-04) 0.0 (2.04e-03) 0.0 (2.14e-03)
Qwen2.5Coder-14B vs Qwen2.5Coder-32B 423.0 (2.83e-27) 100.5 (2.00e-22) 43.5 (9.29e-23)
Qwen2.5-3B vs Qwen2.5-7B 18.0 (2.43e-08) 0.0 (3.43e-05) 0.0 (3.58e-05)
Qwen2.5-7B vs Qwen2.5-14B 960.5 (1.70e-13) 500.5 (1.65e-11) 574.0 (2.62e-10)
Qwen2.5-14B vs Qwen2.5-32B 1925.0 (1.87e-04) 1858.5 (6.60e-03) 2108.5 (1.43e-02)
Table 29: Wilcoxon signed-rank test results comparing GPT model variants. Each cell shows the test statistic (p-value). We use the Bonferroni corrected significance threshold $\alpha=0.025$.
Model Comparison Detection Localization P Localization R
GPT-4o vs GPT-4.1 2550.0 (9.21e-01) 3961.0 (5.83e-01) 2207.5 (4.48e-06)
o3-mini vs o4-mini 575.0 (6.68e-01) 1171.0 (1.44e-03) 1079.0 (1.47e-04)
G.3 Failure Analysis of MetaLint CoT Model VS Non CoT Model
We observe that a significant portion of the lower detection recall of the CoT MetaLint Qwen3-4B model, relative to its non CoT counterpart, can be attributed to its higher tendency to predict NO VIOLATIONS FOUND in cases that do, in fact, contain violations. Specifically, the CoT model fails to flag violations in 89 additional instances compared to the non CoT model, amounting to nearly 17% of the evaluation set (89 out of 536 examples).
The idiom wise distribution of these missed violations is shown in Figure 4. While the failure distribution follows a somewhat long tail pattern, the most significant drops occur for PEP 614, PEP 616, and PEP 593. Notably, if the CoT model matched the non CoT model’s performance on just these three PEPs, its detection recall would rise to 0.605, surpassing that of all open source baselines evaluated.
Upon inspecting CoT traces for these and other idioms (see examples in Table 30), we identify several recurring failure modes: 1) Ambiguity in interpreting the idiom specification. For example, in PEP 614, which targets decorators with complex expressions, the CoT model often labels expressions that humans consider complex as simple. 2) Overthinking and repetitive reasoning traces, particularly for PEP 616. 3) Skipping or entirely missing lines that contain violations, again observed in PEP 616. 4) Underspecified idioms. For instance, in PEP 593, which recommends using the Annotated type from the typing module to attach metadata to type hints, the spec lacks clarity and concrete examples, making it hard to learn what constitutes a violation.
We also find similar issues in idioms like PEP 487, which discourages the use of metaclasses for simple customization tasks that could be handled via __init_subclass__ or __set_name__. The CoT model often misclassifies such ``simple'' use cases as complex.
Overall, these patterns suggest that the CoT model applies the idiom specifications more conservatively, resulting in higher precision but at the cost of reduced recall.
Figure 4: Distribution of comparative failures of the CoT MetaLint Qwen3-4B model relative to its non-CoT variant. While errors span a long tail across many PEPs, the majority are concentrated in three: PEP614, PEP593, and PEP616, which motivates our focused analysis on these cases.
Table 30: Example chains of thought for various PEPs where the CoT model incorrectly flags NO VIOLATIONS FOUND instead of the non CoT model.
PEP Chain of Thought Reason for Failure
614
… I need to check the code file for violations of the 614 idiom,
which is about using complex expressions in decorators. …
… The register variable is defined on line 7 as partial(register,
server=hippa). So, the decorator is a partial function. …
… the idiom allows for complex expressions in the decorator.
For example, if the decorator was something like
@lambda func: timed(…)(func), that would be allowed. …
… the register is a partial function, which is a simple expression. …
… the decorator is just a variable that's a partial function. …
Ambiguity about what constitutes
a complex expression. The idiom
targets decorators with complex
expressions that make code
more verbose
525
… The user … wants me to check for … idiom, 525, which
is about using asynchronous generators with "async for"
syntax … the violation would be if the code uses
synchronous generators (like regular for loops) instead of
asynchronous generators (using async for) … Wait, the
example given in the idiom is a class that uses __aiter__
and __anext__ to create an asynchronous iterator. The fix
is to use a function with async def and yield. So, I need to
check if there are any instances where the code is using
synchronous generators (like for loops) instead of
asynchronous ones. … Looking through the code, I
don't see any for loops that are using 'async for'. …
CoT accidentally looks for the
incorrect construct at a crucial
point (async for instead of for)
487
… using __init_subclass__ and __set_name__ for class
customization without metaclasses … using metaclasses
for simple customization tasks makes code unnecessarily
complex … there's a class called FirestoreClientMeta,
which is a metaclass … used for registration of transports
… the problem is whether this is a simple customization …
not just for a single customization but for managing
multiple transports … the idiom is about using metaclasses
for simple customization … but in this case … handling a
more complex task … maybe this is not a violation …
Ambiguity around what constitutes
a simple class. The idiom targets
instances where metaclasses
are used for simple customization
tasks that could have been handled
with "__init_subclass__" or
"__set_name__")
Table 31: Cross-Idiom Generalization on JEP & PMD Best Practice Idioms: Effect of different MetaLint training setups (SFT and RS-DPO) on Llama3.2-3B-Instruct (Table 31). The transfer column indicates training and test data on the left and right side of the arrow. Best score across the compared training setups per model are bolded.
Model Transfer Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Det}$ $R_{Det}$ $F_{Det}$
Llama3.2-3B-Instruct PMD $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
Llama3.2-3B-Instruct + SFT 0.2251 0.4421 0.2983 0.2822 0.2778 0.28
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4395 0.8908 0.5886 0.593 0.5969 0.5949
Llama3.1-8B-Instruct PMD $\rightarrow$ PMD 0.3656 0.4015 0.3827 0.1253 0.131 0.1281
Llama3.1-8B-Instruct + SFT 0.2264 0.4508 0.3014 0.3201 0.3152 0.3177
Llama3.1-8B-Instruct + SFT + RS-DPO 0.4427 0.9191 0.5976 0.6506 0.6709 0.6606
Llama3.2-3B-Instruct PMD $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
Llama3.2-3B-Instruct + SFT 0.2286 0.4072 0.2928 0.1626 0.1336 0.1467
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4903 0.8338 0.6175 0.4216 0.3333 0.3721
Llama3.1-8B-Instruct PMD $\rightarrow$ JEP 0 0 0 0 0 0
Llama3.1-8B-Instruct + SFT 0.2166 0.3724 0.2739 0.1455 0.1142 0.128
Llama3.1-8B-Instruct + SFT + RS-DPO 0.4964 0.8047 0.614 0.4615 0.3395 0.3912
Llama3.2-3B-Instruct JEP $\rightarrow$ JEP 0.3855 0.0096 0.0187 0.0005 0.0004 0.0005
Llama3.2-3B-Instruct + SFT 0.9567 0.8411 0.8952 0.7837 0.754 0.7686
Llama3.2-3B-Instruct + SFT + RS-DPO 0.9406 0.86 0.8985 0.7859 0.7651 0.7753
Llama3.1-8B-Instruct JEP $\rightarrow$ JEP 0 0 0 0 0 0
Llama3.1-8B-Instruct + SFT 0.9658 0.8466 0.9023 0.809 0.7844 0.7965
Llama3.1-8B-Instruct + SFT + RS-DPO 0.9308 0.8686 0.8986 0.8131 0.7756 0.7939
Llama3.2-3B-Instruct JEP $\rightarrow$ PMD 0.0457 0.0079 0.0134 0.0015 0.0022 0.0017
Llama3.2-3B-Instruct + SFT 0.3722 0.2708 0.3152 0.0574 0.0869 0.0692
Llama3.2-3B-Instruct + SFT + RS-DPO 0.4322 0.4054 0.4183 0.0878 0.1222 0.1022
Llama3.1-8B-Instruct JEP $\rightarrow$ PMD 0.3656 0.4015 0.3827 0.1253 0.131 0.1281
Llama3.1-8B-Instruct + SFT 0.3514 0.2229 0.2728 0.0383 0.0753 0.0508
Llama3.1-8B-Instruct + SFT + RS-DPO 0.436 0.4898 0.4613 0.0831 0.1351 0.1029
Table 32: Egregious Localization Error Types: Breakdown of recurring error categories observed in a manual analysis of 50 egregious localization failures made by the Qwen3-4B non-CoT model trained with SFT+RS-DPO. An egregious localization error is defined as predicting a non-empty set of line numbers that is completely disjoint from the non-empty ground-truth set, indicating correct detection but failed localization.
Error Type Definition Ocurrences (%)
Adjacent Predicts a line immediately adjacent to the true violating line, but not the correct one. 12%
Before/After Flags a line that follows the best practice rather than the line that violates it. 4%
Definition vs. Initialization Flags an initialization instead of the definition that actually constitutes the violation. 4%
Plausible-Looking Flags a superficially plausible but ultimately incorrect line due to surface-level similarity or insufficient semantic nuance. 46%
Random Selects an apparently arbitrary line, suggesting a collapse in localization despite correct detection. 28%
Repetition Repeats the same incorrect line multiple times, indicating a decoding or sampling artifact. 6%